### v1

# Multi-Agent Deep Reinforcement Learning for Zonal Ancillary Market Coupling

Abstract
We characterize zonal ancillary market coupling relying on noncooperative game theory. To that purpose, we formulate the ancillary market as a multi-leader single follower bilevel problem, that we subsequently cast as a generalized Nash game with side constraints and nonconvex feasibility sets. We determine conditions for equilibrium existence and show that the game has a generalized potential game structure. To compute market equilibrium, we rely on two exact approaches: an integrated optimization approach and Gauss-Seidel best-response, that we compare against multi-agent deep reinforcement learning. On real data from Germany and Austria, simulations indicate that multi-agent deep reinforcement learning achieves the smallest convergence rate but requires pretraining, while best-response is the slowest. On the economics side, multi-agent deep reinforcement learning results in smaller market costs compared to the exact methods, but at the cost of higher variability in the profit allocation among stakeholders. Further, stronger coupling between zones tends to reduce costs for larger zones.
Index Terms:
Ancillary Markets, Game Theory, MARL.
I Introduction
A characteristic feature of electricity is the impossibility to store it in large proportions. Hence, it must be produced exactly when it is consumed, which raises a need for a real-time monitoring of the balance between production and consumption to guarantee the stability of the power grid. When imbalances occur, mechanisms are automatically triggered to restore the nominal frequency of the electric system: a change in power production is made by some production units, which are then said to provide ancillary services, in exchange of a monetary contribution. In Western Europe, many market zones are progressively switching from an automatic and mandatory involvement of qualified production units with an amount that is fixed by the system operator to a market mechanism based on daily auctions in which the market operator publishes a demand and then the producers submit offer curves that are aggregated until they meet the demand. From the point of view of a market actor, submitting a bid in such an auction leaves some space for strategic bidding, which requires some information about the market participants. This knowledge, together with the design of a relevant bidding strategy, should enable market actors to engage their production assets in the best way in ancillary services markets, which will then give them a fair idea of the revenue they can expect. Although these ancillary services markets exist in many countries, the bidding zones can be somehow associated: for instance the Austrian and German areas can share their bids, subject to some constraints and to the availability of the transmission network across both countries. Market participants might be willing to investigate the impact of this shared market structure and to see if it stabilizes their revenues over time. The main challenges to address are the characterization of the market equilibria and the understanding of their evolution when market zones get interconnected: from the perspective of the market operator, it is important to gain insight into the behaviors of the stakeholders and to gauge the opportunity to enforce market integration over multiple countries if this can, e.g., mitigate price peak events in one country thanks to interconnection. In this article, we use noncooperative game theory to model ancillary services markets. Market equilibria are computed relying on reformulations of the game and multi-agent deep reinforcement learning (MARL). Our flexible framework allows us to assess the consequences of the interconnection of market zones by dynamically changing the coupling factors. The approach can be generalized to any number of zones. Our main contributions are summarized as follows:
  • we formulate a multi-zone ancillary market model as a bilevel problem, that we subsequently reformulate as a generalized Nash game with generalized potential structure. We obtain theoretical guarantees for the existence of market equilibria;
  • we propose three algorithmic solutions for the market model: two exact approaches to compute an equilibrium, integrated optimization and Gauss-Seidel best-response, and a multi-agent deep reinforcement learning (MARL) simulation. The latter allows us to run a decentralized simulation of the energy market with strategic learning and non-strategic agents;
  • we compare the three algorithm performance on real data from the Germany-Austria market, computing social cost and fairness indicators. Finally, we test multiple coupling settings for the market, observing a decrease in the market operators’ costs with higher couplings.
This paper is organized as follows: in Section I-A, we introduce the relevant literature; in Section II, we formulate the ancillary service market as a noncooperative game. In Section III, we derive conditions for existence of market equilibria, that can be subsequently computed relying on the algorithms introduced in Section IV. Finally, in Section V, we implement the algorithms on real data from Austria and Germany, that we compare numerically using performance measures and economic metrics, concluding in Section VI.
I-A Related Work
There are two main directions for research on ancillary (and zonal) market design: working on the design of the market itself, studying different formulations and their features [1]; or working on the strategic bidding aspect [2], focusing more on the point of view of the energy producers. In [3], the author discusses possible ways of clearing the electricity market and selecting bids, proposing and analyzing different options for the market operator’s objective function, the settlement rules and the pricing of the products. Previous works are mostly theoretical, comparing the different possible market designs relying on noncooperative game theory. The authors of [4] instead, propose a market designed to account for the presence of wind power producers. On the same line is the two parts work of [5, 6], where the authors first discuss the design of an ancillary service focused on primary frequency response, and then simulate this design on different test cases. Another example exploring the impact of renewable based generators, is [7], discussing bidding strategies for ancillary markets with participation of battery energy storage systems. We conclude the discussion on the formulation and features of ancillary markets with [8], which explores the consequences and opportunities of implementing ancillary services at distribution grid level, considering frequency response, voltage regulation and other parameters. Regarding the computing of bidding strategies and solutions for ancillary and zonal markets, a first contribution can be found in [9], where the authors model the market as a two-players game, analyzing the strategies of two producers going one against the other, and solving the game to obtain different market equilibria. In [10] a control based, decentralized solution is proposed to create ancillary services helping with power imbalances deriving from renewable sources. The authors in [11] formulate a zonal market as a two-stage stochastic problem, which they then recast as a mixed-integer bilinear program in order to solve it. This solution is computed from an external point of view, considering all producers at once. In [12], the market is modeled as a Stackelberg game where the market operator is the leader and the producers are the followers; the solution is then obtained by recasting the problem as a mathematical program with equilibrium constraints. Finally, [13] focuses on a game theoretic model with nonlinear convex constraints for the optimization of each producer, which in turn is simulated using a linear dynamic model. The authors then analyze theoretically the existence of a variational equilibrium and propose a distributed algorithm to compute it. The use of learning algorithm for strategic bidding has become more common in the last years, an example is [14], where the authors use deep reinforcement learning to model the bidding strategy of intelligent electricity producers. Regarding ancillary markets, the application of learning based algorithm is not particularly widespread, nonetheless, there are some examples. In [15], the authors focus on ancillary markets with hydro-based producers. The simulation they propose uses discrete inputs for the algorithm and discretized bids for the learning producers, having market power. The results are not definitive, but the learning agents show sign of strategic behavior. A more advanced approach can be found in [16], where the implemented learning algorithm utilizes continuous inputs and outputs and is specifically designed to handle multi-agent settings. The producers here only bid a price, instead of a price-quantity couple as in our case, and the focus is on the grid optimization.
Notations
We denote $\mathbb{I}$ the indicator function. Given $\mathcal{A}\subset\mathbb{R}^{n}$, $|\mathcal{A}|$ refers to the cardinality of the set. Let $x_{n}\in\mathbb{R}^{m_{n}}$, with $m_{n}\in\mathbb{N}^{\star}$, we define $\text{col}((x_{n})_{n})$ as the stack of the $(x_{n})_{n}$ vectors. Given a vector of decision variables $x:=(x)_{n\in\mathcal{N}}$, we define $x_{-n}:=(x_{m})_{m\in\mathcal{N}\setminus\{n\}}$. When we write $\nabla^{2}f(x^{\star})$, with $x^{\star}\in\mathbb{R}^{d}$, $d\in\mathbb{R}^{\star}$, we consider the Hessian matrix of the function $f$ evaluated in $x^{\star}$, where the element $(i,j)$ refers to $\partial_{x_{i},x_{j}}f(x^{\star}),1\leq i,j\leq d$.
TABLE I: List of variables, index and sets.
Symbol Meaning
$\mathcal{N}$ Set of $N$ producers
$\mathcal{Z}$ Set of $Z$ zones
$z(n)$ Zone of producer $n$
$\mathcal{N}(z)$ Subset of producers in $z$
$K_{n}$ Maximum number of bids for producer $n$
$B$ Total number of bids
$\Delta_{n,k}$ Capacity offered by producer $n$ for bid $k$
$\Delta_{n}^{min},\Delta_{n}^{max}$ Lower and upper capacity for producer $n$
$\pi_{n,k}$ Price offered by producer $n$ for bid $k$
$\pi_{n}^{min},\pi_{n}^{max}$ Lower and upper price for producer $n$
$x_{n,k}^{z}$ Fraction of accepted bid $k$ of producer $n$ in zone $z$
II An Adversarial Game Formulation
The ancillary services market can be formulated as a multi-leader single follower bilevel optimization problem. At the upper level, producers (leaders) bid strategically to maximize their revenues while anticipating the market clearing. At the lower level, the market operator (follower) determines the fractions of the bids that are accepted in order to minimize the total cost it has to pay to the producers to meet the zonal demands, while satisfying zonal, interzone and bid-related constraints. Because of the opposite objectives of the producers and market operator, the problem can be interpreted as an adversarial game, also called a zero sum game.
We define $\mathcal{N}:=\{1,\ldots,N\}$ as the set of $N$ producers and $\mathcal{Z}:=\{1,...,Z\}$ as the set of zones. Each producer $n\in\mathcal{N}$ can bid up to $K_{n}\in\mathbb{N}^{\star}$ bids. Thus, the maximum number of bids is $B:=\sum_{n\in\mathcal{N}}K_{n}$. Further, each zone $z\in\mathcal{Z}$ has three input parameters: the total requested demand $D_{z}\geq 0$, the export limit $E_{z}\geq 0$ and the core portion $C_{z}\geq 0$. For all $n\in\mathcal{N}$, $z(n)\in\mathcal{Z}$ defines the zone of producer $n$ and $\mathcal{N}(z)\subseteq\mathcal{N}$ is the subset of producers in zone $z$. Each producer $n$’s decision variable is a finite sequence of bids $y_{n}:=\text{col}(\Delta_{n},\pi_{n})=\text{col}\Big{(}(\Delta_{n,1},\pi_{n,1}),\dots,(\Delta_{n,K_{n}},\pi_{n,K_{n}})\Big{)}$, where $\Delta_{n}$ is producer $n$’s finite sequence of offers containing for each bid, the offered capacity and price. The decision variables of the market operator are $x:=(x_{n,k}^{z})_{n,k,z}\in[0,1]^{B\times Z}$, where $x_{n,k}^{z}$ represents the fraction of bid activated in each zone. A summary of the definitions is in Tab. I.
II-A Producers
Each producer $n\in\mathcal{N}$ has a capacity $\Delta^{max}_{n}\geq 0$ and a marginal price $\pi^{min}_{n}\geq 0$, and can either submit a single price (with full capacity) or a finite sequence of price and quantity pairs satisfying the following constraints involving finite upper and lower bounds:
$\pi^{min}_{n}\leq\pi_{n,k}\leq\pi^{max}_{n},\quad\forall\;k\in\{1,\dots,K_{n}\},$ (1a)
$\Delta^{min}_{n}\leq\Delta_{n,k}\leq\Delta^{max}_{n},\quad\forall\;k\in\{1,\dots,K_{n}\},$ (1b)
$\sum_{k=1}^{K_{n}}\Delta_{n,k}\leq\Delta^{max}_{n}.$ (1c)
[…]
$J_{n}(y_{n},x_{n}):=\sum_{k=1}^{K_{n}}\pi_{n,k}\Delta_{n,k}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\right).$ (2)
II-B Market Operator
The market operator deals with four sets of constraints that depend on the capacities $\Delta_{n,k},\forall k\in[1,K_{n}],n\in\mathcal{N}$ submitted by the producers:
Zonal Demand
$\sum_{n\in\mathcal{N}}\sum_{1\leq k\leq K_{n}}x_{n,k}^{z}\Delta_{n,k}\geq D_{z},\quad\forall z\in\mathcal{Z}.$ (3)
Capacity Limit
$\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\leq 1,\quad\forall k\in[1,\dots,K_{n}],\quad\forall n\in\mathcal{N}.$ (4)
Export Limit
$\sum_{n\in\mathcal{N}(z)}\sum_{1\leq k\leq K_{n}}\sum_{z^{\prime}\neq z}x_{n,k}^{z^{\prime}}\Delta_{n,k}\leq E_{z},\quad\forall z\in\mathcal{Z}.$ (5)
Core Portion
$\sum_{n\in\mathcal{N}(z)}\sum_{1\leq k\leq K_{n}}x_{n,k}^{z}\Delta_{n,k}\geq C_{z},\quad\forall z\in\mathcal{Z}.$ (6)
Eq. (3) ensures that the total demand in each zone is met, while Eq. (5) and Eq. (6) couple the different zones. In particular, Eq. (6) ensures that there is a given amount of the demand of a specific zone satisfied by the producers in that zone, while Eq. (5) gives a cap to the amount of energy that can be exported. Finally, Eq. (4) captures the fact that for a given bid $(n,k)$, the sum of the fractions activated in each zone is normalized. Constraints (3)-(6) can be compactly written under the form
$\Phi(x,y)\leq 0,$
and we denote $\mathcal{X}(y):=\{x\in[0,1]^{B\times Z}:\Phi(x,y)\leq 0\}$ the feasibility set of the market operator which depends on the profile of strategies of the producers $y:=(y_{n})_{n}$.
The market operator aims to minimize the total cost it has to pay to the producers. Its objective function takes the form:
$J_{\text{MO}}(x,y):=\sum_{n\in\mathcal{N}}\sum_{1\leq k\leq K_{n}}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\right)\Delta_{n,k}\pi_{n,k}.$ (7)
The market operator determines the fractions of bids to activate in each zone $(x_{n,k}^{z})_{n,k,z}$, under constraints (3)-(6):
$\min_{x}$ $\quad J_{\text{MO}}(x,y),$ (8a)
s.t. $\quad x\in\mathcal{X}(y).$ (8b)
We report a schematic view of a two-zone market in Fig. 1, where we highlight with arrows the different zonal constraints.
Zone AZone BZonal Demand AZonal Demand BExportCore PortionCore PortionP1P2P3P4P5
Figure 1: Representation of an ancillary market with two zones, with three producers in the first zone and two producers in the second zone. Constraints are highlighted with arrows.
II-C Market Clearing and KKT Reformulation
We formulate the ancillary market clearing as a pessimistic multi-leader single follower bilevel optimization problem:
$\forall n\in\mathcal{N},\quad$ $\min_{x_{n}}\max_{y_{n}}$ $\quad J_{n}(y_{n},x_{n}),$ (9a)
s.t. $x\in\operatorname*{argmin}_{x\in\mathcal{X}(y)}\quad J_{\text{MO}}(x,y),$ (9b)
$y\in\prod_{j}\mathcal{Y}_{j}.$ (9c)
With this formulation, each producer optimizes its bids without considering the different zones. Such a distinction is made by the market operator when it determines the fraction of bids accepted in each zone. Further, each producer’s objective function is independent of the other producers’ strategies. Coupling between the objective functions occurs through the market operator’s strategy. In case of multiple solutions at the lower level, the pessimistic formulation enables the selection of the worst equilibrium for the producers, obtaining a more robust formulation of the market. If the $\min$ operator were to be removed in Eq. (9a), multiple solutions might be considered at the lower level (9b)-(9c).
The Lagrangian function associated with the market clearing defined in problem (8) takes the form:
$L(x,y;\Lambda)=\sum_{n\in\mathcal{N}}\sum_{k=1}^{K_{n}}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\right)\Delta_{n,k}\pi_{n,k}+\Lambda^{\top}\Phi(x,y),$ (10)
where $\Lambda\in\mathbb{R}_{+}^{3Z+B}$ refers to the dual variables associated with constraints (3)-(6). The analytical expression of the Lagrangian function in (10) can be found in Appendix A.
Lemma 1 (Slater’s Constraint Qualification).
Slater’s constraint qualification holds if one of these conditions is satisfied:
  • (i)
    ${\sum_{n\in\mathcal{N}(z)}}\Delta_{n}^{\max}>D_{z},\;\forall z\in\mathcal{Z}$,
  • (ii)
    ${\sum_{z^{\prime}\neq z}}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\Delta_{n}^{\max}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{\max}$
    and $D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{\max}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$.
Proof.
Under condition (i), each zone can satisfy its own demand; thus, there exists an interior-point solution. Under condition (ii), one zone $z\in\mathcal{Z}$ may not be able to meet its own demand, i.e., $D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{\max}>0$. In this situation, we need enough demand offers left from other zones to cover the lack of supply, thus $\sum_{z^{\prime}\neq z}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\Delta_{n}^{\max}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{max}$. Since the export constraint should hold, we get that $D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{\max}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$. ∎
When the lower-level problem (8) satisfies Lemma 1, we can replace it by its KKTs in (9). Hence, each producer has to anticipate the decision $x$ of the market operator, meaning that in the single level the decision variables for producer $n$ will be $y_{n}\in\mathcal{Y}_{n}$ and $x_{n}\in\mathcal{X}_{n}(y)$. Before writing the full reformulation, we introduce tracking variables for every $z\in\mathcal{Z}$:
$d^{z}_{n}(x_{-n},y_{-n}):=D_{z}-\sum_{j\neq n}\sum_{1\leq k\leq K_{j}}x_{j,k}^{z}\Delta_{j,k},$ (11a)
$e^{z}_{n}(x_{-n},y_{-n}):=E_{z}-\sum_{\begin{subarray}{c}j\in\mathcal{N}(z),\\ j\neq n\end{subarray}}\sum_{z^{\prime}\neq z}\sum_{1\leq k\leq K_{j}}x_{j,k}^{z^{\prime}}\Delta_{j,k},$ (11b)
$c^{z}_{n}(x_{-n},y_{-n}):=C_{z}-\sum_{\begin{subarray}{c}j\in\mathcal{N}(z),\\ j\neq n\end{subarray}}\sum_{1\leq k\leq K_{j}}x_{j,k}^{z}\Delta_{j,k}.$ (11c)
The single-level problem for agent $n$ is
$\min_{x_{n}}\,\max_{y_{n},\Lambda_{n}}J_{n}(y_{n},x_{n}),$ (12a)
[…]
$\nabla_{x_{n}}L(x_{n},y_{n};\Lambda_{n})=0,$ (12c)
$\lambda_{n}^{z}\Bigg{(}\sum_{1\leq k\leq K_{n}}x_{n,k}^{z}\Delta_{n,k}-d_{n}^{z}(x_{-n},y_{-n})\Bigg{)}=0,\forall z\in\mathcal{Z},$ (12d)
$\mu_{n,k}\Bigg{(}\sum_{z\in\mathcal{Z}}x_{n,k}^{z}-1\Bigg{)}=0,\;\forall k\in\{1,\dots,K_{n}\},$ (12e)
$\sigma_{n}\Bigg{(}\sum_{k=1}^{K_{n}}\Big{(}\sum_{z^{\prime}\neq z(n)}x_{n,k}^{z^{\prime}}\Big{)}\Delta_{n,k}-e_{n}^{z(n)}(x_{-n},y_{-n})\Bigg{)}=0,$ (12f)
$\delta_{n}\Bigg{(}\sum_{k=1}^{K_{n}}x_{n,k}^{z(n)}\Delta_{n,k}-c_{n}^{z(n)}(x_{-n},y_{-n})\Bigg{)}=0.$ (12g)
Eq. (12c) is the stationarity condition of the lower-level problem (9b)-(9c). Though there is no explicit coupling between the producers’ objective functions, Eqs. (3), (5), (6) and the slackness conditions in Eqs. (12d), (12f), (12g) impose couplings through bids and zones.
We now introduce a more compact notation. First we stress that, given the dependence of the feasibility set $\mathcal{X}$ on $y$, the decision variable $x_{n}$ can be expressed as an implicit function in $y$, leading to $x_{n}\equiv x_{n}(y)$. This in turn, allows us to write the objective function as an explicit function of all producers’ decision variables, i.e. $J_{n}(y_{n},y_{-n})$. Then we let $w_{n}:=(\pi_{n,k},\Delta_{n,k},x_{n,k}^{z},\lambda_{n}^{z},\mu_{n,k},\sigma_{n},\delta_{n})\in\mathcal{W}_{n}:=[\pi^{min}_{n},\pi_{n}^{max}]^{K_{n}}\times[\tilde{\Delta}_{n},\Delta_{n}^{max}]^{K_{n}}\times[0,1]^{K_{n}\cdot Z}\times[0,+\infty]^{Z}\times[0,+\infty]^{K_{n}}\times[0,+\infty]\times[0,+\infty]$. $\mathcal{W}_{n}$ is a polyhedron with dimensionality $T:=3K+Z(K+1)+2$. The set of $l_{n}$ local constraints are captured by $h_{n}:\omega_{n}\rightarrow\mathbb{R}^{l_{n}}$ and that of the $r$ shared constraints by $G:\Omega\rightarrow\mathbb{R}^{r}$, where $\Omega_{n}\supseteq\mathcal{W}_{n}$ and $\Omega:=\prod_{n=1}^{N}\Omega_{n}$. In particular $h_{n}$ and $G$ are continuously differentiable mappings. Finally, we introduce the local feasibility set for each player $n$: $\widetilde{\mathcal{W}}_{n}:=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0\}$.
Using the previously introduced compact notation with the explicit objective function, we get $J_{n}(w_{n},w_{-n})$. Thus, the $N$ optimization problems in Eqs. (12) give rise to a generalized Nash game $\Gamma$ that we denote in strategic form $\Gamma:=(\mathcal{N},\prod_{n}\mathcal{W}_{n},(J_{n})_{n})$:
$\forall n\in\mathcal{N},\quad$ $\max_{w_{n}\in\mathcal{W}_{n}}$ $J_{n}(w_{n},w_{-n}),$ (13a)
s.t. $h_{n}(w_{n})\leq 0,$ (13b)
$G(w_{n},w_{-n})\leq 0.$ (13c)
III Theoretical Framework
In a large part of the literature, the convexity of the agents’ optimization problems is one indispensable assumption under which noncooperative games are studied and analyzed. On the contrary, in this paper, $\Gamma$ is a generalized Nash game with side constraints, involving trilinear objective functions and nonconvex constraints thus giving rise to a nonconvex game. In order to proceed we will use results established in [17], where the authors study a class of problems with similar settings. To that purpose, we first apply the generic setting of [17]. Then, we present a relaxed version of our problem, which is more tractable for this analysis, providing results regarding the equivalence with the non-penalized version. Further, we prove existence of market equilibrium under mild assumptions, and highlight the generalized potential structure of $\Gamma$.
III-A Nonconvex Game with Side Constraints
The formulation of [17] considers a problem where the coupling constraints are treated as penalties with associated prices, in a modified objective function. This problem is then solved as an $(N+1)$-player game, where $N$ players optimize their variables and an external one optimizes the prices, which are shared by all players. In our formulation the coupling constraints are Eqs. (3)-(5)-(6)-(12d)-(12f)-(12g), while the remaining part is considered as local constraints.
With the compact notation previously introduced, we can write the problem as an $(N+1)$-player game with prices [^1] Notice that to directly extend the results of [17], we now work with the negative of the objective function and in minimization settings.:
$\forall n\in\mathcal{N},\quad\min_{w_{n}\in\widetilde{\mathcal{W}}_{n}}\left[-J_{n}(w_{n},w_{-n})+\sum_{j=1}^{r}\rho_{j}G_{j}(w)\right]$ (14)
where we call the externally optimized price, $\rho\in\mathbb{R}^{r}_{+}$. We can now recall the definition of Nash Equilibrium that is a relevant solution concept to analyze this class of games:
Definition 1 (Nash Equilibrium with Side Constraints, [17, Definition 1]).
A Nash equilibrium (NE) for the game (14) is a strategy-price tuple $(w^{\star},\rho^{\star})$, such that for all $n$:
$w_{n}^{\star}\in\operatorname*{argmin}_{w_{n}\in\widetilde{\mathcal{W}}_{n}}\left[-J_{n}(w_{n},w_{-n}^{\star})+\sum_{j=1}^{r}\rho_{j}^{\star}G_{j}(w_{n},w_{-n}^{\star})\right],$
and
$0\leq\rho^{\star}\perp G(w^{\star})\leq 0$
III-B Relaxed Generalized Nash Game
In order to extend the results from [17], we consider a relaxed version of our problem, which makes it more tractable and it will allow us to use mild assumptions to obtain an equilibrium existence guarantee. We add the constraints in Eqs. (3) and (6) to the objective function, which gives the function $J_{n}^{p}(w)$, the new constraint maps $h_{n}^{p}(w_{n})$ and $G^{p}(w)$, and the associated set of penalties $M_{n}$ for each agent (see Appendix B). We can formulate the penalized generalized Nash game $\Gamma^{p}$ as follows:
$\min_{w_{n}\in\mathcal{W}_{n}}$ $-J_{n}^{p}(w)$ (15a)
s.t. $h_{n}^{p}(w_{n})\leq 0$ (15b)
$G^{p}(w_{n},w_{-n})\leq 0$ (15c)
Proposition 1.
If $K\geq\frac{Z}{3}$, there exists a set of penalties $(M_{j})_{j\in\mathcal{N}}$ such that a solution of the penalized game (15) is a solution of $\Gamma$.
Proposition 1 results from the comparisons of the sets of KKTs of the generalized Nash game $\Gamma$ and the penalized game (15) denoted $\Gamma^{p}$. Details can be found in Appendix C.
III-C Existence of a Generalized Nash Equilibrium for $\Gamma$
To obtain existence result for the relaxed problem, we have to verify assumptions on the variables’ domain of definition, the objective functions and the constraint sets.
Assumptions 1.
We assume the following:
  • (a)
    $\mathcal{W}_{n}$ are polyhedral and compact for all $n$;
  • (b)
    $J_{n}(\cdot,w_{-n})$ is continuously differentiable in $w_{n}$, $h_{n}(\cdot)$ and $G(\cdot,w_{-n})$ are twice continuously differentiable in $w_{n}$ for all $n$;
  • (c)
    There exists $w^{\sharp}\in\mathcal{W}:=\prod_{n\in\mathcal{N}}\mathcal{W}_{n}$ such that $h(w_{n}^{\sharp})<0,\,\forall n$ and $G_{n}(w_{n}^{\sharp},w_{-n}^{\sharp})<0,\,\forall n\in\mathcal{N}$;
  • (d)
    The Hessians matrices $\nabla^{2}h_{n}^{p}(w_{n})$ are copositive;
  • (e)
    The Hessians matrices $\nabla^{2}G_{n}^{p}(w)$ are copositive.
Theorem 1 (Existence of a solution for $\Gamma$).
Given $\mathcal{W}$, we assume there exists $\mathcal{W}^{\star}\subseteq\mathcal{W}$, such that for any $w\in\mathcal{W}^{\star}$, the Hessian matrix of the objective function is positive semi-definite. Then, if Assumptions 1 hold, we conclude that inside $\mathcal{W}^{\star}$ the penalized problem has a Nash equilibrium and, consequently, the penalized Generalized Nash game $\Gamma$ also has a Nash equilibrium.
Proof.
Assumption 1 (a) hold for our settings, since the definition of $\mathcal{W}_{n}$ creates polyhedral and compact sets, i.e., we can express any condition of the form $x\in[x_{a},x_{b}]$ as a set of inequalities $x\leq x_{b},x\geq x_{a}$, which are defined as intersections between closed half spaces.
The objective function is continuous and differentiable everywhere in the domain $\mathcal{W}^{\star}$, and the constraints functions are all bilinear or trilinear. Thus, they are all twice continuous and differentiable over the whole domain $\mathcal{W}^{\star}$. Assumption 1 (b) is then satisfied.
Assumption 1 (c) requires the existence of an interior point. Considering $w^{\sharp}$ that strictly satisfies Eq. (3), (4), (5) and (6), we immediately obtain that $\lambda_{n}^{z}=0,\mu_{n,k}=0,\sigma_{n}=0,\delta_{n}=0,\,\forall k\in[1,K_{n}],\,\forall n\in\mathcal{N},\,\forall z\in\mathcal{Z}$. Eq. (12c) needs instead to be relaxed to a strict inequality.
Finally, for Assumptions 1 (d)-(e), we need to compute the Hessians matrices associated to the constraints functions. Regarding the local set of constraints $h_{n}^{p}(w_{n})$, we obtain $\nabla^{2}h_{n}^{p}(w_{n})=0,\,\forall n\in\mathcal{N},\,\forall w_{n}\in\mathcal{W}_{n}$, except for Eq. (12c), for which the Hessian (we leave the full form in Appendix D) is easily obtained and it satisfies the copositivity requirement. Considering the only remaining global constraint, Eq (5), we obtain a constant positive Hessian (full computation in the Appendix), which directly satisfies the copositivity assumption.
We can now apply [17, Theorem 13] to the $(N+1)-$players game to obtain the existence of a Nash equilibrium. Further, using results from [17, Sec. 2.1], we have an equivalence between the Nash equilibrium of the $(N+1)-$players game and the associated generalized Nash game, which in our case is the penalized version $\Gamma^{p}$. We conclude using Proposition 1, which states the equivalence between $\Gamma^{p}$ and $\Gamma$. ∎
We conclude this section checking the potential structure of the game:
Proposition 2.
The generalized Nash game described in Eq. (13) is a generalized potential game (GPG).
Proof.
Using [18, Def. 2.1], we have two requirements for a GNG to be a GPG. The first requirement amounts to have a feasible set that is nonempty, which is easily satisfied in our setting. In particular, the feasible set, which we call $C$, does not have to be the cartesian product of the players’ feasible sets, meaning that $C\subseteq\prod_{n=1}^{N}\bar{\mathcal{W}}_{n}(w_{-n})$, where $\bar{\mathcal{W}}_{n}(w_{-n}):=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0,G(w)\leq 0\}$. The second requirement is for a potential $P(w)$ function to exist, which in our case is satisfied by using $P(w):=\sum_{n=1}^{N}J_{n}(w)$. ∎
From Proposition 2, we can rewrite Eqs. (13) as:
$\max_{w\in W}$ $\quad\sum_{m=1}^{N}J_{n}(w_{m},w_{-m}),$ (16a)
s.t. $\quad h_{n}(w_{n})\leq 0,\forall n,$ (16b)
$\quad G(w_{n},w_{-n})\leq 0,\,\forall n,$ (16c)
which can be solved using a standard nonlinear programming solver. The potential structure is also interesting algorithmically, as it can be used to obtain a convergence guarantee for the best-response algorithm.
IV Algorithmic Solutions
We solve the market problem relying on integrated optimization and Gauss-Seidel best-response method, that we compare against a MARL algorithm. For the integrated optimization approach we use the MINOS solver, which relies on a projected Lagrangian method [19]. Next, we discuss the best-response algorithm we use and how we built the multi-agent simulation.
IV-A Gauss-Seidel Best Response
We follow the implementation of the Gauss-Seidel Best Response algorithm from [18]. Here we remove the coordination between players, and we solve the problem in Eq. (13) for each agent iteratively (once again with the MINOS solver), stopping the whole process once the decision variables of the agents do not change anymore, as summarized in Alg. 1.
Algorithm 1 Gauss-Seidel Best Response.
1:Initialize problem with feasible point $w^{0}=[w_{1}^{0},\dots,w_{N}^{0}]$, $k=0$
2:while Termination criterion is false do
3:     for $n\in[1,\dots,N]$ do
4:         Solve regularized Eq. (13) for agent $n$ with $[w_{1}^{k+1},\dots,\allowbreak w_{n}^{k},\dots,\allowbreak w_{N}^{k+1}]$
5:     end for
6:     Update regularization parameter
7:     $k\leftarrow k+1$
8:end while
The integrated solution can be seen as a coordinated solution for all the producers at once, while using the Best-Response algorithm allows for each producer to optimize its strategy alone. In the results section we will explain more in details how these two solutions differ in modeling the interactions between agents.
Furthermore, from [18, Theorem 5.2] and the potential structure of our game, we have a guarantee of convergence to a Nash equilibrium, provided it exists, which we already proved in Section III-C.
These two approaches allow us to compare fully coordinated and uncoordinated interaction schemes between agents, that we will compare against the multi-agent simulation based on reinforcement learning.
IV-B Multi Agent Deep Reinforcement Learning
The core idea of reinforcement learning is to have an agent learning directly from exposure to an unknown environment. It is formally modeled as a Markov Decision Process (MDP) $\{S,A,r\}$ where $S$ is the state space, $A$ the action space and $r:S\times A\to\mathbb{R}$ the reward function. We define the discounted return over a trajectory of length $L$ as $R(s,a):=\sum_{t=0}^{t=L-1}\gamma^{t}r(s_{t},a_{t})$, where $10<\gamma<$ 1 is the discount factor. An agent interacts with the environment sampling action from its policy $a\sim\mu(s)$, and it aims at maximizing the reward, so the goal is to find the policy that maximizes the discounted return over an infinite horizon. In particular, defining the $Q$-function as:
$Q_{\mu}(s,a)=\mathbb{E}[R(s,a)|\mu,s_{0}=s,a_{0}=a],$
we want to solve the following optimization to obtain the optimal policy $\pi^{*}$:
$Q^{*}(s,a)=Q_{\mu^{*}}(s,a)=\max_{\mu}Q_{\mu}(s,a).$
Generally, we do not have access to analytical form for neither $Q$ nor $\mu$, making the optimization extremely complex to solve. RL algorithms, and in particular deep RL (DRL) algorithms, try to solve this problem learning a policy that approximates the optimal one. DRL employs neural networks to model both the policy $\mu$ and the $Q$-function, so that the resulting algorithm is model free and can be used for continuous or discrete states and actions. We will not explore in depth here the differences between DRL algorithms, but we will provide a quick overview of the one we use.
We implemented the multi-agent deep reinforcement learning algorithm described in [20]: this is called an actor-critic algorithm, where an actor learns the policy and a critic the $Q$-function. The actor and critic are neural networks parameterized by $\theta$ and $\omega$ respectively, which are trained over specific losses. In particular the loss function for the critic is based on the Bellman’s equation:
$\mathcal{L}_{Q}=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[(Q_{\omega}(s,a)-(r+\beta Q_{\omega}(s^{\prime},\mu_{\theta}(s^{\prime}))))^{2}\right],$
where $\mathcal{D}$ is a buffer where we save previously encountered states, the actions that the agent used and the reward it obtained. The actor instead minimizes the following loss:
$\mathcal{L}_{\mu}=-\mathbb{E}_{s\sim\mathcal{D}}\left[Q_{\omega}(s,\mu_{\theta}(s))\right]$
The intuition behind this algorithm is that the critic learns to map future rewards to specific states and actions (which is the $Q$-function), while the actor tries to maximize the $Q$-function itself (represented by the critic output). What we presented here is the single agent implementation, in our multi-agent settings we use a modified version, where each agent has its own actor and critic, but the critics receive as input the actions of all the agents. This was introduced in [20] as a way of stabilizing the learning and improving the results. In particular, given the zonal nature of our environment, we share the actions between the agents of the same zone, but not between all agents. In Alg. 2 we report in pseudo-code the structure for the training of the RL agents.
We leave further details on the implementation in Appendix E.
Algorithm 2 Basic structure of the zonal MARL algorithm.
1:Initialize problem state $s_{0}$
2:for $t\in L$ do
3:     Get action $a_{n}^{t}\leftarrow\mu_{n}(s_{t}),\forall n\in\mathcal{N}$
4:     Clear market and obtain $(r^{t}_{n},s_{t+1})\;\forall n\in\mathcal{N}$
5:     Build zonal action: $a(z)^{t}=[a_{n}^{t}\;\forall n\in\mathcal{N}(z)]$
6:     Store $(s_{t},a(z)^{t},r^{t}_{n},s^{t+1})$ for each agent
7:     for $z\in\mathcal{Z}$ do
8:         for $n\in\mathcal{N}(z)$ do
9:              Sample $(s,a(z),r_{n},s^{\prime})$ from $\mathcal{D}$
10:              Train critic $Q_{n}$ with $\mathcal{L}_{Q}(s,a(z),r_{n},s^{\prime})$
11:              Train actor $\mu_{n}$ with $\mathcal{L}_{\mu}(s,Q_{n})$
12:         end for
13:     end for
14:end for
V Simulation Results
We now discuss the numerical result of the three algorithmic solutions we propose. We test these algorithms on real data from the German and Austrian ancillary service markets, for the period between January 2024 and August 2024 [21]. The maximum length for the bid sequence is 5, and each submitted capacity must be $\geq 5$ MW. As shown in Tab. II, the demand in Germany is much higher and volatile than that in Austria. Furthermore, the export limit is fixed at $80$ MW for both zones and the core portion is $100$ MW for Austria and 00 MW for Germany. The symmetric and low export limit suggests that producers in Germany could have some impact in Austria, while the converse situation is less likely to arise.
TABLE II: Demand data summary.
Zone Min $25$th perc. $75$th perc. Max
Germany $1745$ MW $1898$ MW $1988$ MW $2103$ MW
Austria $200$ MW $200$ MW $200$ MW $225$ MW
In our settings we consider 8 producers, with the parameters described in Tab. III. The parameters are chosen in order to have a competitive environment, making it more difficult to have one or two producers dominating the whole market. We also observed from the real bids submitted, that the total offered capacity should be close to double the demand; our parameters allow for this to happen.
TABLE III: Producers’ parameters.
Producer Zone $\Delta^{max}$ (MW) $\pi^{min}$
0 0 700 7
1 0 700 7
2 1 150 3
3 1 150 3
4 0 650 6
5 0 600 5
6 0 850 8
7 1 350 4
Finally, we test multiple values for the export constraints, in order to study the effect that the two zones have on each other. All simulations were done on a MacBook Pro with an M2 Pro chip and 16G of memory [^2] The code is available at the repository https://github.com/FrancescoMorri/MARL_ancillary_markets.
V-A Convergence of the Multi-Agent RL Simulation
Figure 2: Episodic reward for each RL based agent throughout the training. Each episode represents passing through the full market data. We plot the rewards divided by zone, each point represent the average reward during an episode and the shaded region the standard deviation.
We start by analyzing the convergence and stability of the multi-agent simulation. Our settings are particularly complicated, since the market environment is dependent on all agents decisions. We tested two settings for the MARL algorithm: one with 4 learning based producers, and one with 8. In the first case, the 4 other producers always bid their marginal price. The second case creates a very dynamic environment, since at every step the price of the market is decided based on the actions of 8 independent RL agents. To assess the stability and convergence, we consider the reward of each agent throughout the training. More precisely, we compute the average reward during an episode, this being a full pass through the data. With this data we obtain Fig. 2, where we plot the episodic average reward for each agent divided by zone, considering the case with only RL based producers. Notice that the reward is not the profit, as we mention in the full description of the algorithm in the Appendix. From Fig. 2 we can see that each agent increases its reward during the initial episodes and eventually plateaus to a stable value. These results are particularly interesting, since they show how the algorithm we use is well suited for a medium scale simulation with competitive learning agents, providing results that we can compare with the classical solutions.
V-B MARL Solution and Classical Solutions
We first discuss the theoretical difference between the integrated solution and the Best-Response solution. Starting from the potential problem, we consider the role of the stationarity condition coming from Eq. (10), which for player $n$, bid $k$ and zone $z$ is the following:
$\begin{split}&\nabla_{x_{n,k}^{z}}L(x;\lambda,\mu,\sigma,\delta)=\Delta_{n,k}\pi_{n,k}-\lambda_{z}\Delta_{n,k}+\mu_{n,k}+\\ &+(1-\mathbb{I}_{z=z(n)})\sigma_{z(n)}\Delta_{n,k}-\mathbb{I}_{z=z(n)}\delta_{z(n)}\Delta_{n,k}.\end{split}$ (17)
Notice that every $\mu$ must be always 00, since every other term is multiplied by $\Delta$ and can be grouped. Furthermore, $\delta$ will be 00 for most realistic cases, since the core portion is usually a small fraction of the total demand, meaning that the constraint is always saturated. $\sigma$ on the other hand may be different from 00, which means that $\lambda$ must be different from 00 too, given the signs in the equation. Lastly, we highlight that $\pi$ is indexed by agent and bid, while $\lambda$ by zone, which means that, in most cases, the prices have to be equal to $\lambda$, virtually creating a pay-as-clear market. On the other hand, when using the Best Response algorithm, every agent computes its own $\lambda$, meaning that they are not forced to bid at the same price, and they can actually bid at very high prices with no repercussions. This constraint creates two very different markets, providing two different solutions to analyze.
Regarding the solution obtained from the MARL algorithm, we have less theoretical guarantees, but we can consider it as a more realistic simulation of the market, since every producer has access only to its own information and its only aim is to increase its profit.
We start by analyzing the average cost to run the market. This value is computed by summing the average profits of each producer over the full data available. In particular, for the MARL producers, we consider the cost for the final episode of their training, where we assume they reached stable bidding strategies. The results are reported in Fig. 3. We can immediately see how the two exact algorithms produce higher costs than the learning-based counterpart, which can be explained by the form of the constraints, as detailed above, and by the lack of actual competition. In the integrated solution, the problem is solved by an exogenous external coordinator, so by definition there will be no competition; while for the best response, each producer is free to update its decision variables uncoupled from the others. We can see the average cost for the MARL simulation with only 4 learning producers is very low, but this is also affected by the presence of static producers that always bid at marginal cost, driving the electricity price down.
Figure 3: Average cost of running the market, obtained with different algorithms.
It is interesting to analyze how the profit is shared between producers in the different solutions. We do so by computing the Gini index [22], which, given a set of profits $p_{1}\leq p_{2}\leq\dots\leq p_{n}$, can be obtained as:
$g=\frac{1}{n}\left(n+1-2\left(\frac{\sum_{i=1}^{n}(n+1-i)p_{i}}{\sum_{i=1}^{n}p_{i}}\right)\right).$
The lower $g$ is, the more equal is the sharing of profits. We report the results of this analysis in Tab. IV. We can see that, even if the BR and potential solutions produce higher costs, they also obtain a better overall equality index, with the two MARL based simulations getting the highest index. Once again this can be explained by the fact that the multi-agent simulation is unregulated: each learning agent can bid anything it wants, which creates a situation where few producers can drive others’ profit almost to zero, using their market power. Looking at the index computed by zone, the differences between classical and learning-based algorithm stay the same.
TABLE IV: We report here the Gini index for all simulations: lower is a more equal share of the total profit between producers.
Simulation Overall Germany Austria
Best Response $0.43$ $0.16$ $0.15$
Potential Solution $0.51$ $0.47$ $0.28$
4 RL $0.64$ $0.52$ $0.67$
8 RL $0.77$ $0.69$ $0.33$
We conclude with a comparison of the computational time required by each algorithm to simulate the market for the full dataset. In Tab. V we report the Evaluation and Training time: the two classical algorithms obviously have no training time, while for the MARL algorithm, the latter is what takes the longest. Although, it is important to stress that the MARL algorithm needs to be trained once and then it can be used to evaluate multiple settings without retraining, or eventually with a much shorter training routine. We also highlight that the BR algorithm’s evaluation time is longer than the training of 8 RL agents, due to the generally slow convergence.
TABLE V: Algorithms time for evaluation and training.
Algorithm Evaluation Time Training Time
Best Response $\sim 8$h -
Potential $\sim 10$m -
4 RL $\sim 3$s $\sim 1.5$h
8 RL $\sim 4$s $\sim 2$h
Figure 4: Average cost of running the market with 8 and 4 RL agents, evaluated over different values for the export constraints. Each element on the x-axis corresponds to a couple of export constraints: G stands for Germany and A for Austria, the constraints are in MW.
V-C Effects of Coupling Between Zones
We now analyze how the two zones influence each other in our model. To do so we simulate the market with a range of different values for the export limit $E_{z}$ in Eq. (5). We write the export limit for Germany (Austria) as a fraction of the maximal demand of Austria (Germany): $E_{G}=c_{G}\cdot\max{D_{A}}$, and we explore the ranges $c_{G}\in[0,2]$ and $c_{A}\in[0,0.2]$, so we go from completely uncoupled zone to a situation where the larger zone (Germany) has an export limit which is double the total demand of the smaller zone (Austria).
To quickly evaluate all of these values we use the MARL simulations, in both cases (4 and 8 RL agents). We report in Fig. 4 the average cost to run the market, computed in the same way as for Fig. 3, for the range of constraints values. In particular, we trained the agents for the two cases with $c_{G}=0.4,c_{A}=0.04$, and used them on all other values. When analyzing these results, the actual costs are not particularly relevant, as they depend on the demand and other external factors, while the trend of increase/decrease is what we have to interpret.
We can see that in the case of $4$ RL agents, the average cost is not influenced by the export constraint: this is likely due to the fact that the presence of static agents, always bidding at marginal price, already drives the cost down, and it is not possible to decrease it further. The case with $8$ RL agents is much more interesting: we observe that the cost for Austria is rather stable, while the one for Germany is basically linearly decreasing. We can interpret this behavior considering the differences between the producers in Austria and Germany: in the smaller zone we have less producers, with lower capacities and marginal prices, while in the larger Germany the producers have higher marginal prices. Once the Austrian producers can enter the German market with their leftover capacity, they can bid the same way they would bid for their local market, which in turns, forces the German producers to bid lower, in order to be more competitive. Furthermore, since it is the market that decides how to divide each bid between zone, a producer can not create a priori a specific bid for a specific zone, each bid needs to be general enough to be competitive on every zone available.
We conclude by mentioning [23], where the authors simulated a more economically sophisticated model of the European electricity market, to test what the addition of another smaller zone would do to the costs in the French zone. Their results show the same trend we observe here, with the smaller zone obtaining similar cost, while the larger one decreases its own. Our main methodological advantage, using a MARL based simulation, is that we can quickly test a range of different settings, not just a single scenario, providing a less detailed but broader view of the effect of different market designs. Further, using pre-trained agents, we are testing how a bidding strategy developed for a certain market design translates to a different one, when there is no time to adapt. This approach is of particular interest for situations where the coupling may change due to grid failures, with no warning.
VI Conclusions
In this work we presented a model for multi-zonal ancillary markets based on a bilevel formulation. Each electricity producer is represented by an optimization problem with local and coupling constraints, with the objective of maximizing its profit, while the market operator objective is to minimize the total cost, so we create an adversarial formulation. We then reformulated the bilevel problem as a single level optimization, from the perspective of each producer, using the KKT reformulation of the market optimization. We interpret the ensemble of single level problems of each producer as a generalized Nash game. The game we obtain has side constraints, trilinear objective functions and nonconvex constraints, making it particularly challenging to solve. We analyze this game considering a relaxed version, and we obtain existence results for the Nash equilibrium that we relate to the initial generalized Nash game. We conclude our theoretical analysis showing that the generalized Nash game has a generalized potential structure.
We then consider different algorithms specifically tailored to our problem: as baselines, we rely on an integrated optimization reformulation relying on the potential structure of the game, and Gauss-Seidel Best Response to solve the generalized Nash game. Further, we implement a multi-agent deep reinforcement learning algorithm to simulate the market in a more realistic, distributed way.
We conclude validating the three algorithmic solutions on real data from the ancillary market between Germany and Austria. Our results showcase that computationally, the MARL simulation is the fastest to evaluate the data, but the agents have to be trained once before, while the best response approach is the slowest. The two classical algorithms also produce solutions with market costs much higher than the multi-agent counterpart, but the latter produces a more unfair split of the profits between the players. Finally, we test different forms of coupling between the zones, showing that a stronger coupling could lead to a decrease in cost for the larger zone, a result which is validated by other analysis present in the literature [23].
Appendix A KKT Reformulation Details
The full form of the Lagrangian for the market clearing problem of Eq. (8) is the following:
$\begin{split}L(x;&\lambda,\mu,\sigma,\delta)=\sum_{n\in\mathcal{N}}\sum_{k=1}^{K_{n}}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\right)\Delta_{n,k}\pi_{n,k}\\ +&\sum_{z\in\mathcal{Z}}\lambda_{z}\left(D_{z}-\sum_{n\in\mathcal{N}}\sum_{k=1}^{K_{n}}x_{n,k}^{z}\Delta_{n,k}\right)\\ +&\sum_{n\in\mathcal{N}}\sum_{k=1}^{K_{n}}\mu_{n,k}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}-1\right)\\ +&\sum_{z\in\mathcal{Z}}\sigma_{z}\left(\sum_{n\in\mathcal{N}(z)}\sum_{k=1}^{K_{n}}\left(\sum_{z^{\prime}\neq z}x_{n,k}^{z^{\prime}}\right)\Delta_{n,k}-E_{z}\right)\\ +&\sum_{z\in\mathcal{Z}}\delta_{z}\left(C_{z}-\sum_{n\in\mathcal{N}(z)}x_{n,k}^{z}\Delta_{n,k}\right),\end{split}$ (18)
where we explicit the dual variables $\lambda\in\mathbb{R}_{\geq 0}^{Z},\mu\in\mathbb{R}_{\geq 0}^{B},\sigma\in\mathbb{R}_{\geq 0}^{Z},\delta\in\mathbb{R}_{\geq 0}^{Z}$.
Appendix B Penalized Objective Function
We report here the explicit form for the penalized objective function used in the penalized GNG of Eqs. (15). As described in Section III-B, we relax constraints Eqs. (3)-(6), to obtain:
$J_{n}^{p}(w)$ $=\sum_{1\leq k\leq K}\sum_{z\in\mathcal{Z}}\pi_{n,k}\Delta_{n,k}x_{n,k}^{z}$ (19a)
$+\sum_{z\in\mathcal{Z}}M_{D_{z}}\left(\sum_{n\in\mathcal{N}}\sum_{1\leq k\leq K}x_{n,k}^{z}\Delta_{n,k}-D_{z}\right)$ (19b)
$+M_{C}\left(\sum_{j\in z(n)}\sum_{0\leq k\leq K}x_{j,k}^{z(n)}\Delta_{j,k}-C_{z(n)}\right).$ (19c)
Appendix C Proof of Proposition 1
We call the two Lagrangian functions associated to the problems $\mathcal{L}_{n}^{P}$ and $\mathcal{L}_{n}^{U}$, and we consider the solution for $U$: $(w_{n}^{*},\Theta_{n}^{*})\in\mbox{SOL}\{\mbox{KKT of }U\}$, where $\Theta$ is the set of Lagrangian multipliers. This solution must respect all the constraints and be a solution of $\nabla\mathcal{L}^{U}(w^{*}_{n},\Theta^{*}_{n})=0$. We now call the part of Lagrangian multipliers that are still present in $P$ $\Theta[\Sigma_{n}]$, and we consider the set of equations coming from $\nabla\mathcal{L}^{P}(w_{n},\Theta_{n}[\Sigma_{n}];M_{n})=0$, where $M_{n}$ is the set of penalties for agent $n$. If we evaluate these equations in $(w_{n}^{*},\Theta_{n}^{*}[\Sigma_{n}])$, we can solve them w.r.t. the penalties $M_{n}$, obtaining $\bar{M}_{n}$. Given the assumptions, solving the system $\nabla\mathcal{L}^{P}(w_{n},\Theta_{n}[\Sigma];\bar{M}_{n})=0$ in $(w_{n},\Theta_{n}[\Sigma])$ provides $\tilde{w}_{n}$ which is a solution also for $\nabla\mathcal{L}^{U}(w_{n},\Theta_{n})=0$.
Appendix D Hessians Analysis
For the stationarity condition of Eq. (12c) we obtain the following cases:
$x_{n}^{\top}\nabla_{x_{n}}^{2}\left(\nabla_{x_{n,k}^{z}}L\right)x_{n}=\begin{cases}\pi_{n,k}-\lambda_{n}^{z}-\delta_{n}\geq 0,\quad\mbox{if }z=z(n)\\ \pi_{n,k}-\lambda_{n}^{z}+\sigma_{n}\geq 0,\quad\mbox{if }z\neq z(n)\end{cases}$
Given that an interior point forces $\lambda_{n}^{z}=0,\delta_{n}=0$ and $\sigma_{n}=0$, and that the prices $\pi_{n,k}$ are positive by definition, both inequalities are always satisfied.
Regarding the Hessian coming from Eq. (5), which we call $\nabla^{2}G_{D_{z}}$, we can observe that deriving twice we get only constant terms, and given that all the terms in the sum are positive, we obtain a Hessian matrix with only $1$s. Then, since our decision variables are all positive by definition, the requirement $x^{\top}\nabla^{2}G_{D_{z}}x\geq 0$ is always true.
Appendix E MADDPG Implementation Details
There are three main problems related to the use of RL algorithms: defining the input state and the output action, mapping the action to satisfy the constraints and deciding the reward function.
State and Action
We choose the following forms for the input state and output action, for each producer:
  • state: $s^{t}_{z}=[p^{t-1},d^{t-1},w^{t-1},\sigma^{t-1}_{z}]$ where $p^{t-1}$ are the previous day clearing prices, $d^{t-1}$ is the day, $w^{t-1}$ is the day of the week and $\sigma_{z}^{t-1}$ is a random signal which is zone-specific.
  • action: $a^{t}_{i}=\left[\left[\pi_{n}^{t,0},\dots,\pi_{n}^{t,K_{n}}\right],\left[\Delta_{n}^{t,0},\dots,\Delta_{i}^{t,K_{n}}\right]\right]$, where $K_{n}$ is the number of bids each producer can submit.
Action Mapping
In our market model we have the following features:
  • Capacity Output: we want each capacity bid to be at least 5 MW and less than the maximum capacity, with also the sum being less or equal than the total capacity of the producer.
  • Price Output: each price bid must be larger than the marginal price $\pi^{min}_{n}$ and smaller than the maximum bid $\pi^{max}_{n}$
We tackle the two outputs separately, building a module for the capacity and one for the price. We use the following mapping:
  • Capacity Module: The output of the capacity module is a vector of length equal to the length of the bid sequence, which is passed through the soft-max function, representing the percentage of capacity to offer at each step. The benefit of this mapping is that is rather simple, with the drawback of forcing each producer to bid its whole capacity at each bid.
  • Price Module: The price module produces two values $\alpha,\beta\geq 0$, which are used to obtain the lowest and highest price of the bidding sequence as: $\pi_{0}=\min(\pi^{max},\pi^{min}+\min(\alpha,\beta))$ and $\pi_{K}=\min(\pi^{max},\pi^{min}+\max(\alpha,\beta))$. We then compute the other prices interpolating between $\pi_{0}$ and $\pi_{K}$.
Reward Design
The most direct approach for the reward function in the market environment would be to use the profit, but it does not encode important information: a producer may place a non-optimal bid, but if the demand is high enough, it will be accepted anyway; further the zones have different demands, so a producer in a certain zone may have a smaller profit just because the demand is lower. What we propose is a reward formed by two terms: a normalized zone profit and a penalty for bid prices that are higher than the clearing price, weighted by the hyperparameter $\beta$:
$r^{t}_{n}=\sum_{z\in\mathcal{Z}}\left(\frac{p_{n,z}^{t}}{d^{t}_{z}/|\mathcal{N}(z)|}-\beta\sum_{k=1}^{K_{n}}\mathbb{I}[\lambda_{z}^{t}>\pi_{n}^{min}]\cdot(\pi_{n}^{t,k}-\lambda_{z}^{t})\right)$ (20)

### v2

@@
-Multi-Agent Deep Reinforcement Learning for Zonal Ancillary Market Coupling
+Nonconvex Game and Multi Agent Reinforcement Learning for Zonal Ancillary Markets
 Abstract
@@
 Our main contributions are summarized as follows:
-  • we formulate a multi-zone ancillary market model as a bilevel problem, that we subsequently reformulate as a generalized Nash game with generalized potential structure.
+  • we formulate a multi-zone ancillary market model as a Stackelberg game, that we subsequently reformulate as a generalized Nash game with generalized potential structure.
 We obtain theoretical guarantees for the existence of market equilibria;
-  • we propose three algorithmic solutions for the market model: two exact approaches to compute an equilibrium, integrated optimization and Gauss-Seidel best-response, and a multi-agent deep reinforcement learning (MARL) simulation.
-The latter allows us to run a decentralized simulation of the energy market with strategic learning and non-strategic agents;
+  • we propose three algorithms to compute market outcomes: an integrated optimization approach and Gauss-Seidel best-response as exact methods, to compute an equilibrium ; a multi-agent deep reinforcement learning (MARL) approach, involving strategic learning and non-learning agents, that can be implemented in a decentralized fashion, reaching a stationary solution;
   • we compare the three algorithm performance on real data from the Germany-Austria market, computing social cost and fairness indicators.
@@
 There are two main directions for research on ancillary (and zonal) market design: working on the design of the market itself, studying different formulations and their features [1]; or working on the strategic bidding aspect [2], focusing more on the point of view of the energy producers.
-In [3], the author discusses possible ways of clearing the electricity market and selecting bids, proposing and analyzing different options for the market operator’s objective function, the settlement rules and the pricing of the products.
-Previous works are mostly theoretical, comparing the different possible market designs relying on noncooperative game theory.
-The authors of [4] instead, propose a market designed to account for the presence of wind power producers.
-On the same line is the two parts work of [5, 6], where the authors first discuss the design of an ancillary service focused on primary frequency response, and then simulate this design on different test cases.
-Another example exploring the impact of renewable based generators, is [7], discussing bidding strategies for ancillary markets with participation of battery energy storage systems.
-We conclude the discussion on the formulation and features of ancillary markets with [8], which explores the consequences and opportunities of implementing ancillary services at distribution grid level, considering frequency response, voltage regulation and other parameters.
-Regarding the computing of bidding strategies and solutions for ancillary and zonal markets, a first contribution can be found in [9], where the authors model the market as a two-players game, analyzing the strategies of two producers going one against the other, and solving the game to obtain different market equilibria.
+In [3], the author discusses possible ways of clearing the electricity market and selecting bids, proposing different options for the market operator’s objective function, the settlement rules and the pricing of the products.
+Previous works are mostly theoretical, comparing the efficiency of different market designs relying on noncooperative game theory.
+In [4] a two-stage electricity market design is considered, where the first stage is the day-ahead market and the second is the real time one.
+The authors obtain proof of equilibrium existence for the model and suggest a market mechanism to reach it.
+On ancillary markets, the two parts work of [5, 6], first discusses the design of an ancillary service focused on primary frequency response, and then simulates this design on different test cases.
+Another example exploring the impact of renewable based generators is [7], discussing bidding strategies for ancillary markets with participation of battery energy storage systems.
+We conclude this part on the formulation and features of ancillary markets with [8], which explores the consequences and opportunities of implementing ancillary services at distribution grid level, considering frequency response, voltage regulation and other parameters.
+Regarding studies of strategic bidding in electricity market, we start by mentioning [9], where the authors focus on the spot market, obtaining a Bayesian Nash equilibrium for the problem, computed in a centralized way and also considering the effect of wind based generation.
+In [4] a two-stage electricity market Regarding the computing of bidding strategies and solutions for ancillary and zonal markets, a first contribution can be found in [10], where the authors model the market as a two-players game, analyzing the strategies of two producers going one against the other, and solving the game to obtain different market equilibria.
 In [11] a control based, decentralized solution is proposed to create ancillary services helping with power imbalances deriving from renewable sources.
@@
 In [13], the market is modeled as a Stackelberg game where the market operator is the leader and the producers are the followers; the solution is then obtained by recasting the problem as a mathematical program with equilibrium constraints.
-Finally, [13] focuses on a game theoretic model with nonlinear convex constraints for the optimization of each producer, which in turn is simulated using a linear dynamic model.
-The authors then analyze theoretically the existence of a variational equilibrium and propose a distributed algorithm to compute it.
-The use of learning algorithm for strategic bidding has become more common in the last years, an example is [14], where the authors use deep reinforcement learning to model the bidding strategy of intelligent electricity producers.
-Regarding ancillary markets, the application of learning based algorithm is not particularly widespread, nonetheless, there are some examples.
+The use of reinforcement learning (RL) as a technique to solve complex problems related to power systems and electricity markets has become more common in the last years, even though applying to real world scenarios remains complicated due to the few theoretical results.
+One interesting work in this direction is [14], where the authors consider constraints to obtain guarantees for the application of an RL algorithm to voltage control.
+Regarding RL for strategic bidding in the electricity markets we mention [15], where the authors use deep reinforcement learning to model the bidding strategy of intelligent electricity producers.
+Focusing specifically ancillary markets, the application of learning based algorithm is not yet widespread, nonetheless, there are some examples.
 In [16], the authors focus on ancillary markets with hydro-based producers.
@@
 $z(n)$ Zone of producer $n$
-$\mathcal{N}(z)$ Subset of producers in $z$
+$\mathcal{N}(z)$ Subset of producers in zone $z$
 $K_{n}$ Maximum number of bids for producer $n$
@@
 II An Adversarial Game Formulation
-The ancillary services market can be formulated as a multi-leader single follower bilevel optimization problem.
+The ancillary services market can be formulated as a multi-leader single follower bilevel problem.
 At the upper level, producers (leaders) bid strategically to maximize their revenues while anticipating the market clearing.
@@
 For all $n\in\mathcal{N}$, $z(n)\in\mathcal{Z}$ defines the zone of producer $n$ and $\mathcal{N}(z)\subseteq\mathcal{N}$ is the subset of producers in zone $z$.
-Each producer $n$’s decision variable is a finite sequence of bids $y_{n}:=\text{col}(\Delta_{n},\pi_{n})=\text{col}\Big{(}(\Delta_{n,1},\pi_{n,1}),\dots,(\Delta_{n,K_{n}},\pi_{n,K_{n}})\Big{)}$, where $\Delta_{n}$ is producer $n$’s finite sequence of offers containing for each bid, the offered capacity and price.
+Each producer $n$’s decision variable is a finite sequence of bids $y_{n}:=\text{col}(\Delta_{n},\pi_{n})=\text{col}\Big{(}(\Delta_{n,1},\pi_{n,1}),\dots,(\Delta_{n,K_{n}},\pi_{n,K_{n}})\Big{)}$, where $\Delta_{n},\pi_{n}$ are producer $n$’s finite sequences of offered capacities and prices, respectively.
 The decision variables of the market operator are $x:=(x_{n,k}^{z})_{n,k,z}\in[0,1]^{B\times Z}$, where $x_{n,k}^{z}$ represents the fraction of bid activated in each zone.
@@
 Eq.
-(3) ensures that the total demand in each zone is met, while Eq.
-(5) and Eq.
-(6) couple the different zones.
-In particular, Eq.
+(3) ensures that the total demand in each zone is met, Eq.
 (6) ensures that there is a given amount of the demand of a specific zone satisfied by the producers in that zone, while Eq.
@@
 and we denote $\mathcal{X}(y):=\{x\in[0,1]^{B\times Z}:\Phi(x,y)\leq 0\}$ the feasibility set of the market operator which depends on the profile of strategies of the producers $y:=(y_{n})_{n}$.
-The market operator aims to minimize the total cost it has to pay to the producers. Its objective function takes the form:
+The market operator aims to minimize the total cost it has to pay to the producers.
+Its objective function takes the form:
 $J_{\text{MO}}(x,y):=\sum_{n\in\mathcal{N}}\sum_{1\leq k\leq K_{n}}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\right)\Delta_{n,k}\pi_{n,k}.$ (7)
@@
 We report a schematic view of a two-zone market in Fig. 1, where we highlight with arrows the different zonal constraints.
-Zone AZone BZonal Demand AZonal Demand BExportCore PortionCore PortionP1P2P3P4P5
 Figure 1: Representation of an ancillary market with two zones, with three producers in the first zone and two producers in the second zone. Constraints are highlighted with arrows.
 II-C Market Clearing and KKT Reformulation
-We formulate the ancillary market clearing as a pessimistic multi-leader single follower bilevel optimization problem:
+We formulate the ancillary market clearing problem as a pessimistic multi-leader single follower bilevel optimization problem:
 $\forall n\in\mathcal{N},\quad$ $\min_{x_{n}}\max_{y_{n}}$ $\quad J_{n}(y_{n},x_{n}),$ (9a)
@@
 In case of multiple solutions at the lower level, the pessimistic formulation enables the selection of the worst equilibrium for the producers, obtaining a more robust formulation of the market.
-If the $\min$ operator were to be removed in Eq.
-(9a), multiple solutions might be considered at the lower level (9b)-(9c).
 The Lagrangian function associated with the market clearing defined in problem (8) takes the form:
@@
 Lemma 1 (Slater’s Constraint Qualification).
-Slater’s constraint qualification holds if one of these conditions is satisfied:
+Slater’s constraint qualification holds if one of these conditions is satisfied
   • (i)
@@
 Then we let $w_{n}:=(\pi_{n,k},\Delta_{n,k},x_{n,k}^{z},\lambda_{n}^{z},\mu_{n,k},\sigma_{n},\delta_{n})\in\mathcal{W}_{n}:=[\pi^{min}_{n},\pi_{n}^{max}]^{K_{n}}\times[\tilde{\Delta}_{n},\Delta_{n}^{max}]^{K_{n}}\times[0,1]^{K_{n}\cdot Z}\times[0,+\infty]^{Z}\times[0,+\infty]^{K_{n}}\times[0,+\infty]\times[0,+\infty]$.
-$\mathcal{W}_{n}$ is a polyhedron with dimensionality $T:=3K+Z(K+1)+2$.
+$\mathcal{W}_{n}$ is a polyhedron with dimensionality $23K_{n}+Z(K_{n}+1)+$ 2.
 The set of $l_{n}$ local constraints are captured by $h_{n}:\omega_{n}\rightarrow\mathbb{R}^{l_{n}}$ and that of the $r$ shared constraints by $G:\Omega\rightarrow\mathbb{R}^{r}$, where $\Omega_{n}\supseteq\mathcal{W}_{n}$ and $\Omega:=\prod_{n=1}^{N}\Omega_{n}$.
@@
 III Theoretical Framework
-In a large part of the literature, the convexity of the agents’ optimization problems is one indispensable assumption under which noncooperative games are studied and analyzed.
+In a large part of the literature, the convexity of the agents’ optimization problems is an essential assumption under which noncooperative games are studied and analyzed.
 On the contrary, in this paper, $\Gamma$ is a generalized Nash game with side constraints, involving trilinear objective functions and nonconvex constraints thus giving rise to a nonconvex game.
-In order to proceed we will use results established in [17], where the authors study a class of problems with similar settings.
-To that purpose, we first apply the generic setting of [17].
-Then, we present a relaxed version of our problem, which is more tractable for this analysis, providing results regarding the equivalence with the non-penalized version.
-Further, we prove existence of market equilibrium under mild assumptions, and highlight the generalized potential structure of $\Gamma$.
-III-A Nonconvex Game with Side Constraints
+In order to proceed we will rely on results established in [18], where the authors study a class of problems with similar settings.
+To that purpose, we first introduce a penalized version of our problem, which is more tractable for this analysis, providing results regarding the equivalence with the non-penalized version; then we present the settings of [18], which we use to prove existence of market equilibrium under mild assumptions.
+Finally we highlight the potential structure of $\Gamma$.
+III-A Penalized Generalized Nash Game
+We consider a penalized version of our problem, which makes it more tractable and it will allow us to use mild assumptions to obtain an equilibrium existence guarantee.
+We add the constraints in Eqs.
+(3) and (6) to the objective function, which gives the function $J_{n}^{p}(w)$ and the new constraint maps $h_{n}^{p}(w_{n})$ and $G^{p}(w)$ (see Appendix B for more details).
+We can formulate the penalized generalized Nash game $\Gamma^{p}$ as follows:
+$\forall n\in\mathcal{N},\quad$ $\min_{w_{n}\in\mathcal{W}_{n}}$ $-J_{n}^{p}(w)$ (14a)
+s.t. $h_{n}^{p}(w_{n})\leq 0$ (14b)
+$G^{p}(w_{n},w_{-n})\leq 0$ (14c)
+Proposition 1.
+If $K_{n}\geq\frac{Z}{3}$, there exists a set of penalties such that a solution of the penalized game (14) is a solution of $\Gamma$.
+Proposition 1 results from the comparisons of the sets of KKTs of the generalized Nash game $\Gamma$ and the penalized game (14) denoted $\Gamma^{p}$. Details can be found in Appendix C.
+Thanks to this reformulation we obtain copositive Hessians matrices associated to the constraints $h_{n}^{p}(w_{n})$ and $G^{p}(w)$, which is the key point in obtaining an existence result.
+We discuss the form of these matrices in Appendix D.
+III-B Nonconvex Game with Side Constraints
 The formulation of [18] considers a problem where the coupling constraints are treated as penalties with associated prices, in a modified objective function.
 This problem is then solved as an $(N+1)$-player game, where $N$ players optimize their variables and an external one optimizes the prices, which are shared by all players.
-In our formulation the coupling constraints are Eqs.
-(3)-(5)-(6)-(12d)-(12f)-(12g), while the remaining part is considered as local constraints.
-With the compact notation previously introduced, we can write the problem as an $(N+1)$-player game with prices [^1] Notice that to directly extend the results of [17], we now work with the negative of the objective function and in minimization settings.:
-$\forall n\in\mathcal{N},\quad\min_{w_{n}\in\widetilde{\mathcal{W}}_{n}}\left[-J_{n}(w_{n},w_{-n})+\sum_{j=1}^{r}\rho_{j}G_{j}(w)\right]$ (14)
-where we call the externally optimized price, $\rho\in\mathbb{R}^{r}_{+}$. We can now recall the definition of Nash Equilibrium that is a relevant solution concept to analyze this class of games:
-Definition 1 (Nash Equilibrium with Side Constraints, [17, Definition 1]).
+With the compact notation previously introduced, we can write the problem in Eq.
+(14) as an $(N+1)$-player game with prices [^1] Notice that to directly extend the results of [18], we now work with the negative of the objective function and in minimization settings.:
+$\forall n\in\mathcal{N},\quad\min_{w_{n}\in\widetilde{\mathcal{W}}^{p}_{n}}\left[-J_{n}^{p}(w_{n},w_{-n})+\sum_{j=1}^{r}\rho_{j}G_{j}^{p}(w)\right]$ (15)
+where we call the externally optimized price, $\rho\in\mathbb{R}^{r}_{+}$ and introduce the penalized local feasibility set $\widetilde{\mathcal{W}}_{n}^{p}:=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}^{p}(w_{n})\leq 0\}$.
+We can now recall the definition of Nash Equilibrium [18, Definition 1] that is a relevant solution concept to analyze this class of games:
+Definition 1 (Nash Equilibrium with Side Constraints).
 A Nash equilibrium (NE) for the game (15) is a strategy-price tuple $(w^{\star},\rho^{\star})$, such that for all $n$:
@@
 $0\leq\rho^{\star}\perp G(w^{\star})\leq 0$
-III-B Relaxed Generalized Nash Game
-In order to extend the results from [17], we consider a relaxed version of our problem, which makes it more tractable and it will allow us to use mild assumptions to obtain an equilibrium existence guarantee.
-We add the constraints in Eqs.
-(3) and (6) to the objective function, which gives the function $J_{n}^{p}(w)$, the new constraint maps $h_{n}^{p}(w_{n})$ and $G^{p}(w)$, and the associated set of penalties $M_{n}$ for each agent (see Appendix B).
-We can formulate the penalized generalized Nash game $\Gamma^{p}$ as follows:
-$\min_{w_{n}\in\mathcal{W}_{n}}$ $-J_{n}^{p}(w)$ (15a)
-s.t. $h_{n}^{p}(w_{n})\leq 0$ (15b)
-$G^{p}(w_{n},w_{-n})\leq 0$ (15c)
-Proposition 1.
-If $K\geq\frac{Z}{3}$, there exists a set of penalties $(M_{j})_{j\in\mathcal{N}}$ such that a solution of the penalized game (15) is a solution of $\Gamma$.
-Proposition 1 results from the comparisons of the sets of KKTs of the generalized Nash game $\Gamma$ and the penalized game (15) denoted $\Gamma^{p}$. Details can be found in Appendix C.
 III-C Existence of a Generalized Nash Equilibrium for $\Gamma$
-To obtain existence result for the relaxed problem, we have to verify assumptions on the variables’ domain of definition, the objective functions and the constraint sets.
+To obtain existence result for the penlaized problem, we have to verify assumptions on the variables’ domain, the objective functions and the constraint sets.
+Thanks to our penalized reformulation, we only need to verify a set of mild assumptions:
 Assumptions 1.
@@
   • (c)
-    There exists $w^{\sharp}\in\mathcal{W}:=\prod_{n\in\mathcal{N}}\mathcal{W}_{n}$ such that $h(w_{n}^{\sharp})<0,\,\forall n$ and $G_{n}(w_{n}^{\sharp},w_{-n}^{\sharp})<0,\,\forall n\in\mathcal{N}$;
-  • (d)
-    The Hessians matrices $\nabla^{2}h_{n}^{p}(w_{n})$ are copositive;
-  • (e)
-    The Hessians matrices $\nabla^{2}G_{n}^{p}(w)$ are copositive.
+    There exists $w^{\sharp}\in\mathcal{W}:=\prod_{n\in\mathcal{N}}\mathcal{W}_{n}$ such that $h_{n}(w_{n}^{\sharp})<0,\,\forall n\in\mathcal{N}$ and $G(w^{\sharp})<0$;
 Theorem 1 (Existence of a solution for $\Gamma$).
-Given $\mathcal{W}$, we assume there exists $\mathcal{W}^{\star}\subseteq\mathcal{W}$, such that for any $w\in\mathcal{W}^{\star}$, the Hessian matrix of the objective function is positive semi-definite.
-Then, if Assumptions 1 hold, we conclude that inside $\mathcal{W}^{\star}$ the penalized problem has a Nash equilibrium and, consequently, the penalized Generalized Nash game $\Gamma$ also has a Nash equilibrium.
+Define $\mathcal{W}^{\star}\subseteq\mathcal{W}$ such that for any $w\in\mathcal{W}^{\star}$, the Hessian matrix of the objective function is positive semi-definite.
+Then, there exists a generalized Nash equilibrium on $\mathcal{W}^{\star}$ solution of $\Gamma^{p}$.
 Proof.
-Assumption 1 (a) hold for our settings, since the definition of $\mathcal{W}_{n}$ creates polyhedral and compact sets, i.e., we can express any condition of the form $x\in[x_{a},x_{b}]$ as a set of inequalities $x\leq x_{b},x\geq x_{a}$, which are defined as intersections between closed half spaces.
-The objective function is continuous and differentiable everywhere in the domain $\mathcal{W}^{\star}$, and the constraints functions are all bilinear or trilinear.
-Thus, they are all twice continuous and differentiable over the whole domain $\mathcal{W}^{\star}$.
+Assumption 1 (a) holds, since $\mathcal{W}_{n}$ is defined as a polyhedral and compact set, i.e., we can express any condition of the form $x\in[x_{a},x_{b}]$ as a set of inequalities $x\leq x_{b},x\geq x_{a}$, which are defined as intersections between closed half spaces.
+Further, the objective function is continuous and differentiable everywhere in the domain $\mathcal{W}^{\star}$, and the constraints functions are bilinear or trilinear.
+Thus, they are twice continuous and differentiable over the whole domain $\mathcal{W}^{\star}$.
 Assumption 1 (b) is then satisfied.
@@
 (12c) needs instead to be relaxed to a strict inequality.
-Finally, for Assumptions 1 (d)-(e), we need to compute the Hessians matrices associated to the constraints functions.
-Regarding the local set of constraints $h_{n}^{p}(w_{n})$, we obtain $\nabla^{2}h_{n}^{p}(w_{n})=0,\,\forall n\in\mathcal{N},\,\forall w_{n}\in\mathcal{W}_{n}$, except for Eq.
-(12c), for which the Hessian (we leave the full form in Appendix D) is easily obtained and it satisfies the copositivity requirement.
-Considering the only remaining global constraint, Eq (5), we obtain a constant positive Hessian (full computation in the Appendix), which directly satisfies the copositivity assumption.
-We can now apply [17, Theorem 13] to the $(N+1)-$players game to obtain the existence of a Nash equilibrium.
-Further, using results from [17, Sec.
+We then apply [18, Theorem 13] to the $(N+1)-$players game to obtain the existence of a generalized Nash equilibrium for $\Gamma^{p}$.
+∎
+Corollary 1.
+The generalized Nash game $\Gamma$ defined in (13) admits a solution.
+Proof.
+Using results from [18, Sec.
 2.1], we have an equivalence between the Nash equilibrium of the $(N+1)-$players game and the associated generalized Nash game, which in our case is the penalized version $\Gamma^{p}$.
@@
 Proof.
-Using [18, Def.
-2.1], we have two requirements for a GNG to be a GPG.
+Using [19, Def.
+2.1], two requirements need to be satisfied for a GNG to be a GPG.
 The first requirement amounts to have a feasible set that is nonempty, which is easily satisfied in our setting.
@@
 which can be solved using a standard nonlinear programming solver.
-The potential structure is also interesting algorithmically, as it can be used to obtain a convergence guarantee for the best-response algorithm.
 IV Algorithmic Solutions
@@
 In the results section we will explain more in details how these two solutions differ in modeling the interactions between agents.
-Furthermore, from [18, Theorem 5.2] and the potential structure of our game, we have a guarantee of convergence to a Nash equilibrium, provided it exists, which we already proved in Section III-C.
+Furthermore, from [19, Theorem 5.2] and the potential structure of our game, we have a guarantee of convergence to a Nash equilibrium, provided it exists, which we already proved in Section III-C.
 These two approaches allow us to compare fully coordinated and uncoordinated interaction schemes between agents, that we will compare against the multi-agent simulation based on reinforcement learning.
@@
 The intuition behind this algorithm is that the critic learns to map future rewards to specific states and actions (which is the $Q$-function), while the actor tries to maximize the $Q$-function itself (represented by the critic output).
-What we presented here is the single agent implementation, in our multi-agent settings we use a modified version, where each agent has its own actor and critic, but the critics receive as input the actions of all the agents.
+In the multi-agent settings we appply a modified version of the single agent implementation presented here, where each agent has its own actor and critic, but the critics receive as input the actions of all the agents.
 This was introduced in [21] as a way of stabilizing the learning and improving the results.
@@
 2 we report in pseudo-code the structure for the training of the RL agents.
-We leave further details on the implementation in Appendix E.
+We leave further details on the implementation and references to the code in Appendix E.
 Algorithm 2 Basic structure of the zonal MARL algorithm.
@@
 V Simulation Results
-We now discuss the numerical result of the three algorithmic solutions we propose.
+We now discuss the numerical results of the three algorithmic solutions we propose.
 We test these algorithms on real data from the German and Austrian ancillary service markets, for the period between January 2024 and August 2024 [22].
@@
 Finally, we test multiple values for the export constraints, in order to study the effect that the two zones have on each other.
-All simulations were done on a MacBook Pro with an M2 Pro chip and 16G of memory [^2] The code is available at the repository https://github.com/FrancescoMorri/MARL_ancillary_markets.
+All simulations were done on a MacBook Pro with an M2 Pro chip and 16G of memory.
 V-A Convergence of the Multi-Agent RL Simulation
@@
 2 we can see that each agent increases its reward during the initial episodes and eventually plateaus to a stable value.
-These results are particularly interesting, since they show how the algorithm we use is well suited for a medium scale simulation with competitive learning agents, providing results that we can compare with the classical solutions.
+These results are particularly interesting, since they show how the algorithm is well suited for a medium scale simulation with competitive learning agents, providing results that we can compare with the classical solutions.
+We want to stress here that, using the MARL algorithm, we compute a stationary solution for the market, without a guarantee of it being an equilibrium.
 V-B MARL Solution and Classical Solutions
@@
 This constraint creates two very different markets, providing two different solutions to analyze.
-Regarding the solution obtained from the MARL algorithm, we have less theoretical guarantees, but we can consider it as a more realistic simulation of the market, since every producer has access only to its own information and its only aim is to increase its profit.
+Regarding the solution obtained from the MARL algorithm, even if there are less theoretical guarantees, we can consider it as a more realistic simulation of the market, since every producer has access only to its own information and its only aim is to increase its profit.
 We start by analyzing the average cost to run the market.
@@
 The game we obtain has side constraints, trilinear objective functions and nonconvex constraints, making it particularly challenging to solve.
-We analyze this game considering a relaxed version, and we obtain existence results for the Nash equilibrium that we relate to the initial generalized Nash game.
+We analyze this game considering a penlaized version, and we obtain existence results for the Nash equilibrium that we relate to the initial generalized Nash game.
 We conclude our theoretical analysis showing that the generalized Nash game has a generalized potential structure.
@@
 Appendix B Penalized Objective Function
-We report here the explicit form for the penalized objective function used in the penalized GNG of Eqs. (15). As described in Section III-B, we relax constraints Eqs. (3)-(6), to obtain:
+We report here the explicit form for the penalized objective function used in the penalized GNG of Eqs. (14). As described in Section III-A, we relax constraints Eqs. (3)-(6), to obtain:
 $J_{n}^{p}(w)$ $=\sum_{1\leq k\leq K}\sum_{z\in\mathcal{Z}}\pi_{n,k}\Delta_{n,k}x_{n,k}^{z}$ (19a)

### v3

@@
-Nonconvex Game and Multi Agent Reinforcement Learning for Zonal Ancillary Markets
+Game Theory and Multi-Agent Reinforcement Learning for Zonal Ancillary Markets
 Abstract
@@
 $\Delta_{n,k}$ Capacity offered by producer $n$ for bid $k$
-$\Delta_{n}^{min},\Delta_{n}^{max}$ Lower and upper capacity for producer $n$
+$\underline{\Delta}_{n},\bar{\Delta}_{n}$ Lower and upper capacity for producer $n$
 $\pi_{n,k}$ Price offered by producer $n$ for bid $k$
-$\pi_{n}^{min},\pi_{n}^{max}$ Lower and upper price for producer $n$
+$\underline{\pi}_{n},\bar{\pi}_{n}$ Lower and upper price for producer $n$
 $x_{n,k}^{z}$ Fraction of accepted bid $k$ of producer $n$ in zone $z$
@@
 II-A Producers
-Each producer $n\in\mathcal{N}$ has a capacity $\Delta^{max}_{n}\geq 0$ and a marginal price $\pi^{min}_{n}\geq 0$, and can either submit a single price (with full capacity) or a finite sequence of price and quantity pairs satisfying the following constraints involving finite upper and lower bounds:
-$\pi^{min}_{n}\leq\pi_{n,k}\leq\pi^{max}_{n},\quad\forall\;k\in\{1,\dots,K_{n}\},$ (1a)
-$\Delta^{min}_{n}\leq\Delta_{n,k}\leq\Delta^{max}_{n},\quad\forall\;k\in\{1,\dots,K_{n}\},$ (1b)
-$\sum_{k=1}^{K_{n}}\Delta_{n,k}\leq\Delta^{max}_{n}.$ (1c)
+Each producer $n\in\mathcal{N}$ has a capacity $\bar{\Delta}_{n}\geq 0$ and a marginal price $\underline{\pi}_{n}\geq 0$, and can either submit a single price (with full capacity) or a finite sequence of price and quantity pairs satisfying the following constraints involving finite upper and lower bounds:
+$\underline{\pi}_{n}\leq\pi_{n,k}\leq\bar{\pi}_{n},\quad\forall\;k\in\{1,\dots,K_{n}\},$ (1a)
+$\underline{\Delta}_{n}\leq\Delta_{n,k}\leq\bar{\Delta}_{n},\quad\forall\;k\in\{1,\dots,K_{n}\},$ (1b)
+$\sum_{k=1}^{K_{n}}\Delta_{n,k}\leq\bar{\Delta}_{n}.$ (1c)
 […]
@@
   • (i)
-    ${\sum_{n\in\mathcal{N}(z)}}\Delta_{n}^{\max}>D_{z},\;\forall z\in\mathcal{Z}$,
+    ${\sum_{n\in\mathcal{N}(z)}}\bar{\Delta}_{n}>D_{z},\;\forall z\in\mathcal{Z}$,
   • (ii)
-    ${\sum_{z^{\prime}\neq z}}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\Delta_{n}^{\max}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{\max}$
-    and $D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{\max}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$.
+    ${\sum_{z^{\prime}\neq z}}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\bar{\Delta}_{n}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}$
+    and $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$.
 Proof.
 Under condition (i), each zone can satisfy its own demand; thus, there exists an interior-point solution.
-Under condition (ii), one zone $z\in\mathcal{Z}$ may not be able to meet its own demand, i.e., $D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{\max}>0$.
-In this situation, we need enough demand offers left from other zones to cover the lack of supply, thus $\sum_{z^{\prime}\neq z}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\Delta_{n}^{\max}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{max}$.
-Since the export constraint should hold, we get that $D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{\max}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$.
+Under condition (ii), one zone $z\in\mathcal{Z}$ may not be able to meet its own demand, i.e., $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}>0$.
+In this situation, we need enough demand offers left from other zones to cover the lack of supply, thus $\sum_{z^{\prime}\neq z}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\bar{\Delta}_{n}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}$.
+Since the export constraint should hold, we get that $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$.
 ∎
@@
 $\min_{x_{n}}\,\max_{y_{n},\Lambda_{n}}J_{n}(y_{n},x_{n}),$ (12a)
-[…]
+$\mbox{s.t. }y_{n}\in\mathcal{Y}_{n},x_{n}\in\mathcal{X}_{n}(y),\Lambda_{n}\geq 0,$ (12b)
 $\nabla_{x_{n}}L(x_{n},y_{n};\Lambda_{n})=0,$ (12c)
@@
 $J_{n}(y_{n},y_{-n})$.
-Then we let $w_{n}:=(\pi_{n,k},\Delta_{n,k},x_{n,k}^{z},\lambda_{n}^{z},\mu_{n,k},\sigma_{n},\delta_{n})\in\mathcal{W}_{n}:=[\pi^{min}_{n},\pi_{n}^{max}]^{K_{n}}\times[\tilde{\Delta}_{n},\Delta_{n}^{max}]^{K_{n}}\times[0,1]^{K_{n}\cdot Z}\times[0,+\infty]^{Z}\times[0,+\infty]^{K_{n}}\times[0,+\infty]\times[0,+\infty]$.
-$\mathcal{W}_{n}$ is a polyhedron with dimensionality $23K_{n}+Z(K_{n}+1)+$ 2.
+Then we let $w_{n}:=(\pi_{n,k},\Delta_{n,k},x_{n,k}^{z},\lambda_{n}^{z},\mu_{n,k},\sigma_{n},\delta_{n})\in\mathcal{W}_{n}:=[\underline{\pi}_{n},\bar{\pi}_{n}]^{K_{n}}\times[\tilde{\Delta}_{n},\bar{\Delta}_{n}]^{K_{n}}\times[0,1]^{K_{n}\cdot Z}\times[0,+\infty]^{Z}\times[0,+\infty]^{K_{n}}\times[0,+\infty]\times[0,+\infty]$.
+$\mathcal{W}_{n}$ is a polyhedron with dimensionality $3K_{n}+Z(K_{n}+1)+2$.
 The set of $l_{n}$ local constraints are captured by $h_{n}:\omega_{n}\rightarrow\mathbb{R}^{l_{n}}$ and that of the $r$ shared constraints by $G:\Omega\rightarrow\mathbb{R}^{r}$, where $\Omega_{n}\supseteq\mathcal{W}_{n}$ and $\Omega:=\prod_{n=1}^{N}\Omega_{n}$.
@@
 and
-$0\leq\rho^{\star}\perp G(w^{\star})\leq 0$
+$0\leq\rho^{\star}\perp G(w^{\star})\leq 0.$
 III-C Existence of a Generalized Nash Equilibrium for $\Gamma$
@@
 3:  for $n\in[1,\dots,N]$ do
-4:         Solve regularized Eq. (13) for agent $n$ with $[w_{1}^{k+1},\dots,\allowbreak w_{n}^{k},\dots,\allowbreak w_{N}^{k+1}]$
+4:   Solve regularized Eq. (13) for agent $n$ with $w_{1}^{k+1},\dots,\allowbreak w_{n}^{k},\dots,\allowbreak w_{N}^{k+1}]$
 5:  end for
@@
 It is formally modeled as a Markov Decision Process (MDP) $\{S,A,r\}$ where $S$ is the state space, $A$ the action space and $r:S\times A\to\mathbb{R}$ the reward function.
-We define the discounted return over a trajectory of length $L$ as $R(s,a):=\sum_{t=0}^{t=L-1}\gamma^{t}r(s_{t},a_{t})$, where $10<\gamma<$ 1 is the discount factor.
+We define the discounted return over a trajectory of length $L$ as $R(s,a):=\sum_{t=0}^{t=L-1}\gamma^{t}r(s_{t},a_{t})$, where $0<\gamma<1$ is the discount factor.
 An agent interacts with the environment sampling action from its policy $a\sim\mu(s)$, and it aims at maximizing the reward, so the goal is to find the policy that maximizes the discounted return over an infinite horizon.
@@
 II, the demand in Germany is much higher and volatile than that in Austria.
-Furthermore, the export limit is fixed at $80$ MW for both zones and the core portion is $100$ MW for Austria and 00 MW for Germany.
+Furthermore, the export limit is fixed at $80$ MW for both zones and the core portion is $100$ MW for Austria and 0 MW for Germany.
 The symmetric and low export limit suggests that producers in Germany could have some impact in Austria, while the converse situation is less likely to arise.
@@
 TABLE III: Producers’ parameters.
-Producer Zone $\Delta^{max}$ (MW) $\pi^{min}$
+Producer Zone $\bar{\Delta}$ (MW) $\underline{\pi}$
 0 0 700 7
@@
 $\begin{split}&\nabla_{x_{n,k}^{z}}L(x;\lambda,\mu,\sigma,\delta)=\Delta_{n,k}\pi_{n,k}-\lambda_{z}\Delta_{n,k}+\mu_{n,k}+\\ &+(1-\mathbb{I}_{z=z(n)})\sigma_{z(n)}\Delta_{n,k}-\mathbb{I}_{z=z(n)}\delta_{z(n)}\Delta_{n,k}.\end{split}$ (17)
-Notice that every $\mu$ must be always 00, since every other term is multiplied by $\Delta$ and can be grouped.
-Furthermore, $\delta$ will be 00 for most realistic cases, since the core portion is usually a small fraction of the total demand, meaning that the constraint is always saturated.
-$\sigma$ on the other hand may be different from 00, which means that $\lambda$ must be different from 00 too, given the signs in the equation.
+Notice that every $\mu$ must be always 0, since every other term is multiplied by $\Delta$ and can be grouped.
+Furthermore, $\delta$ will be 0 for most realistic cases, since the core portion is usually a small fraction of the total demand, meaning that the constraint is always saturated.
+$\sigma$ on the other hand may be different from 0, which means that $\lambda$ must be different from 0 too, given the signs in the equation.
 Lastly, we highlight that $\pi$ is indexed by agent and bid, while $\lambda$ by zone, which means that, in most cases, the prices have to be equal to $\lambda$, virtually creating a pay-as-clear market.
@@
   • Capacity Output: we want each capacity bid to be at least 5 MW and less than the maximum capacity, with also the sum being less or equal than the total capacity of the producer.
-  • Price Output: each price bid must be larger than the marginal price $\pi^{min}_{n}$ and smaller than the maximum bid $\pi^{max}_{n}$
+  • Price Output: each price bid must be larger than the marginal price $\underline{\pi}_{n}$ and smaller than the maximum bid $\bar{\pi}_{n}$
 We tackle the two outputs separately, building a module for the capacity and one for the price. We use the following mapping:
@@
 The benefit of this mapping is that is rather simple, with the drawback of forcing each producer to bid its whole capacity at each bid.
-  • Price Module: The price module produces two values $\alpha,\beta\geq 0$, which are used to obtain the lowest and highest price of the bidding sequence as: $\pi_{0}=\min(\pi^{max},\pi^{min}+\min(\alpha,\beta))$ and $\pi_{K}=\min(\pi^{max},\pi^{min}+\max(\alpha,\beta))$.
+  • Price Module: The price module produces two values $\alpha,\beta\geq 0$, which are used to obtain the lowest and highest price of the bidding sequence as: $\pi_{0}=\min(\bar{\pi},\underline{\pi}+\min(\alpha,\beta))$ and $\pi_{K}=\min(\bar{\pi},\underline{\pi}+\max(\alpha,\beta))$.
 We then compute the other prices interpolating between $\pi_{0}$ and $\pi_{K}$.
@@
 What we propose is a reward formed by two terms: a normalized zone profit and a penalty for bid prices that are higher than the clearing price, weighted by the hyperparameter $\beta$:
-$r^{t}_{n}=\sum_{z\in\mathcal{Z}}\left(\frac{p_{n,z}^{t}}{d^{t}_{z}/|\mathcal{N}(z)|}-\beta\sum_{k=1}^{K_{n}}\mathbb{I}[\lambda_{z}^{t}>\pi_{n}^{min}]\cdot(\pi_{n}^{t,k}-\lambda_{z}^{t})\right)$ (20)
+$r^{t}_{n}=\sum_{z\in\mathcal{Z}}\left(\frac{p_{n,z}^{t}}{d^{t}_{z}/|\mathcal{N}(z)|}-\beta\sum_{k=1}^{K_{n}}\mathbb{I}[\lambda_{z}^{t}>\underline{\pi}_{n}]\cdot(\pi_{n}^{t,k}-\lambda_{z}^{t})\right)$ (20)

### v4

@@
 We characterize zonal ancillary market coupling relying on noncooperative game theory.
-To that purpose, we formulate the ancillary market as a multi-leader single follower bilevel problem, that we subsequently cast as a generalized Nash game with side constraints and nonconvex feasibility sets.
-We determine conditions for equilibrium existence and show that the game has a generalized potential game structure.
-To compute market equilibrium, we rely on two exact approaches: an integrated optimization approach and Gauss-Seidel best-response, that we compare against multi-agent deep reinforcement learning.
-On real data from Germany and Austria, simulations indicate that multi-agent deep reinforcement learning achieves the smallest convergence rate but requires pretraining, while best-response is the slowest.
-On the economics side, multi-agent deep reinforcement learning results in smaller market costs compared to the exact methods, but at the cost of higher variability in the profit allocation among stakeholders.
-Further, stronger coupling between zones tends to reduce costs for larger zones.
-Index Terms:
-Ancillary Markets, Game Theory, MARL.
+We formulate the ancillary market as a multi-leader single follower Stackelberg game, that we subsequently cast as a generalized Nash game with side constraints and nonconvex feasible sets.
+We determine conditions for the existence of an equilibrium and show that the game has a generalized potential structure.
+To compute market equilibrium, we rely on two exact approaches: an integrated optimization reformulation involving complementarity constraints, that can be solved using a nonlinear optimization solver, and Gauss-Seidel best-response.
+We compare these exact methods against multi-agent deep reinforcement learning (MARL).
+Simulations on real data from Germany and Austria ancillary markets show that MARL better capture intricate interactions between market agents under varying information-sharing scenarios, achieving faster learning rates to reach a stationary solution, than both exact approaches.
+Further, MARL results in smaller market costs compared to the exact methods, at the cost of higher variability in the profit allocation among stakeholders.
+Finally, on the policy side, we show numerically that tighter coupling between adjacent supply zones reduces the average market cost of the largest zone.
 I Introduction
@@
 In Western Europe, many market zones are progressively switching from an automatic and mandatory involvement of qualified production units with an amount that is fixed by the system operator to a market mechanism based on daily auctions in which the market operator publishes a demand and then the producers submit offer curves that are aggregated until they meet the demand.
-From the point of view of a market actor, submitting a bid in such an auction leaves some space for strategic bidding, which requires some information about the market participants.
-This knowledge, together with the design of a relevant bidding strategy, should enable market actors to engage their production assets in the best way in ancillary services markets, which will then give them a fair idea of the revenue they can expect.
+From the point of view of a market actor, submitting a bid in such an auction leaves some space for strategic bidding, which requires information about the other market participants.
+This information, together with the design of a relevant bidding strategy, should enable market actors to engage their production assets in the best way in ancillary services markets, which will then give them a fair idea of the revenue they can expect.
 Although these ancillary services markets exist in many countries, the bidding zones can be somehow associated: for instance the Austrian and German areas can share their bids, subject to some constraints and to the availability of the transmission network across both countries.
@@
 The approach can be generalized to any number of zones.
+I-A Related Work
+There are two main directions for research on ancillary (and zonal) market design: a) working on the design of the market itself, studying different formulations and their features [lecadre2019]; b) working on the strategic bidding aspect [morri2024], focusing more on the point of view of the energy producers.
+In [oren_design_2001], Oren discusses possible ways of clearing the electricity market and selecting bids, proposing different options for the market operator’s objective function, the settlement rules and the pricing of the products.
+Previous works are mostly theoretical, comparing the efficiency of different market designs relying on noncooperative game theory.
+A new market design is proposed by Ela et al.
+in [ela_effective_2012], taking into account the uncertainty resulting from the massive penetration of wind power in the grid.
+In the same vein, [ela_market_2014, ela_market_2014_2] discuss the design of an ancillary services market focused on primary frequency response, and then simulate this design under different scenarios.
+Another example exploring the impact of renewable based generators is [merten2020bidding], which tackles the optimization of bidding strategies for ancillary markets involving battery energy storage systems.
+We conclude the discussion on the formulation and features of ancillary services markets with [oureilidis_ancillary_2020], which explores the pros.
+and cons.
+of ancillary services market implementation at the distribution grid level, considering frequency response, voltage power regulation and other grid related parameters.
+Regarding the computation of bidding strategies and solutions for ancillary and zonal markets, a first contribution can be found in [morinec2008optimal], where the authors model the market as a two-player noncooperative game, analyzing the strategies of the two producers, and solving the game to obtain different market equilibria.
+In [meyn_ancillary_2014], a control based decentralized solution is proposed to create ancillary services helping with power imbalances caused by renewable generation sources.
+Sarfati and Holmberg in [sarfati_simulation_2020] formulate a zonal market as a two-stage stochastic problem, which they then recast as a mixed-integer bilinear program that they subsequently solve relying on a nonlinear optimization solver.
+This solution is computed from an external point of view, considering all producers at once.
+In [jay2020game], the market is modeled as a Stackelberg game where the market operator is the leader and the producers are the followers; the solution is then obtained by recasting the problem as a mathematical program with equilibrium constraints.
+Finally, [deng_distributed_2024] focuses on a game-theoretic model with nonlinear convex feasible sets, which in turn is solved by simulation relying on a linear dynamic model.
+The authors then analyze theoretically the existence of a variational equilibrium and propose a distributed algorithm to compute it.
+The use of learning algorithm for strategic bidding has become more common in the last years.
+In [ye_deep_2020], a day-ahead electricity market bidding problem with multiple strategic producers is formulated as a Markov game with unobservable information, which is solved relying on a model-free and data-driven approach.
+Regarding ancillary markets, the application of learning based algorithm is not particularly widespread, nonetheless, there are some examples.
+In [abgottspon_strategic_2013], Abgottspon and Andersson focus on ancillary markets with hydro-based producers.
+Their simulation approach uses discrete inputs and discretized bids for the learning producers.
+The results are not definitive, but the learning agents show sign of strategic behavior.
+A more advanced approach can be found in [du_approximating_2021], where the implemented learning algorithm utilizes continuous inputs and outputs and is specifically designed to handle multi-agent settings.
+The producers here only bid a price, instead of a price-quantity couple as in our case, and the focus is on the grid optimization.
+I-B Contributions and Organization
 Our main contributions are summarized as follows:
@@
 We obtain theoretical guarantees for the existence of market equilibria;
-  • we propose three algorithms to compute market outcomes: an integrated optimization approach and Gauss-Seidel best-response as exact methods, to compute an equilibrium ; a multi-agent deep reinforcement learning (MARL) approach, involving strategic learning and non-learning agents, that can be implemented in a decentralized fashion, reaching a stationary solution;
-  • we compare the three algorithm performance on real data from the Germany-Austria market, computing social cost and fairness indicators.
-Finally, we test multiple coupling settings for the market, observing a decrease in the market operators’ costs with higher couplings.
-This paper is organized as follows: in Section I-A, we introduce the relevant literature; in Section II, we formulate the ancillary service market as a noncooperative game.
-In Section III, we derive conditions for existence of market equilibria, that can be subsequently computed relying on the algorithms introduced in Section IV.
-Finally, in Section V, we implement the algorithms on real data from Austria and Germany, that we compare numerically using performance measures and economic metrics, concluding in Section VI.
-I-A Related Work
-There are two main directions for research on ancillary (and zonal) market design: working on the design of the market itself, studying different formulations and their features [1]; or working on the strategic bidding aspect [2], focusing more on the point of view of the energy producers.
-In [3], the author discusses possible ways of clearing the electricity market and selecting bids, proposing different options for the market operator’s objective function, the settlement rules and the pricing of the products.
-Previous works are mostly theoretical, comparing the efficiency of different market designs relying on noncooperative game theory.
-In [4] a two-stage electricity market design is considered, where the first stage is the day-ahead market and the second is the real time one.
-The authors obtain proof of equilibrium existence for the model and suggest a market mechanism to reach it.
-On ancillary markets, the two parts work of [5, 6], first discusses the design of an ancillary service focused on primary frequency response, and then simulates this design on different test cases.
-Another example exploring the impact of renewable based generators is [7], discussing bidding strategies for ancillary markets with participation of battery energy storage systems.
-We conclude this part on the formulation and features of ancillary markets with [8], which explores the consequences and opportunities of implementing ancillary services at distribution grid level, considering frequency response, voltage regulation and other parameters.
-Regarding studies of strategic bidding in electricity market, we start by mentioning [9], where the authors focus on the spot market, obtaining a Bayesian Nash equilibrium for the problem, computed in a centralized way and also considering the effect of wind based generation.
-In [4] a two-stage electricity market Regarding the computing of bidding strategies and solutions for ancillary and zonal markets, a first contribution can be found in [10], where the authors model the market as a two-players game, analyzing the strategies of two producers going one against the other, and solving the game to obtain different market equilibria.
-In [11] a control based, decentralized solution is proposed to create ancillary services helping with power imbalances deriving from renewable sources.
-The authors in [12] formulate a zonal market as a two-stage stochastic problem, which they then recast as a mixed-integer bilinear program in order to solve it.
-This solution is computed from an external point of view, considering all producers at once.
-In [13], the market is modeled as a Stackelberg game where the market operator is the leader and the producers are the followers; the solution is then obtained by recasting the problem as a mathematical program with equilibrium constraints.
-The use of reinforcement learning (RL) as a technique to solve complex problems related to power systems and electricity markets has become more common in the last years, even though applying to real world scenarios remains complicated due to the few theoretical results.
-One interesting work in this direction is [14], where the authors consider constraints to obtain guarantees for the application of an RL algorithm to voltage control.
-Regarding RL for strategic bidding in the electricity markets we mention [15], where the authors use deep reinforcement learning to model the bidding strategy of intelligent electricity producers.
-Focusing specifically ancillary markets, the application of learning based algorithm is not yet widespread, nonetheless, there are some examples.
-In [16], the authors focus on ancillary markets with hydro-based producers.
-The simulation they propose uses discrete inputs for the algorithm and discretized bids for the learning producers, having market power.
-The results are not definitive, but the learning agents show sign of strategic behavior.
-A more advanced approach can be found in [17], where the implemented learning algorithm utilizes continuous inputs and outputs and is specifically designed to handle multi-agent settings.
-The producers here only bid a price, instead of a price-quantity couple as in our case, and the focus is on the grid optimization.
+  • we propose three algorithms to compute market outcomes: an integrated optimization approach and Gauss-Seidel best-response as exact methods; a multi-agent deep reinforcement learning (MARL) approach, that can be implemented in a decentralized fashion, reaching a stationary solution;
+  • we compare the three algorithm performance on real data from the Germany-Austria market, comparing social cost, producers’ profit variability and computation times.
+Finally, comparing multiple coupling scenarios, we observe a decrease in the largest zone market cost with tighter couplings.
+This paper is organized as follows: in Section II, we formulate the ancillary services market as a noncooperative game, in Section III, we derive conditions for existence of market equilibria, that can be subsequently computed relying on the algorithms introduced in Section IV.
+Finally, in Section V, we implement the algorithms on real data from Austria and Germany, that we compare numerically using key performance metrics.
+We conclude in Section VI.
 Notations
-We denote $\mathbb{I}$ the indicator function.
 Given $\mathcal{A}\subset\mathbb{R}^{n}$, $|\mathcal{A}|$ refers to the cardinality of the set.
 Let $x_{n}\in\mathbb{R}^{m_{n}}$, with $m_{n}\in\mathbb{N}^{\star}$, we define $\text{col}((x_{n})_{n})$ as the stack of the $(x_{n})_{n}$ vectors.
-Given a vector of decision variables $x:=(x)_{n\in\mathcal{N}}$, we define $x_{-n}:=(x_{m})_{m\in\mathcal{N}\setminus\{n\}}$.
-When we write $\nabla^{2}f(x^{\star})$, with $x^{\star}\in\mathbb{R}^{d}$, $d\in\mathbb{R}^{\star}$, we consider the Hessian matrix of the function $f$ evaluated in $x^{\star}$, where the element $(i,j)$ refers to $\partial_{x_{i},x_{j}}f(x^{\star}),1\leq i,j\leq d$.
+Given a vector of decision variables $x:=(x_{n})_{n\in\mathcal{N}}$, we define $x_{-n}:=(x_{m})_{m\in\mathcal{N}\setminus\{n\}}$ as the concatenation of all the players’ decision variables except $n$.
+When we write $\nabla^{2}f(x^{\star})$, with $x^{\star}\in\mathbb{R}^{d}$, $d\in\mathbb{R}^{\star}$, we consider the Hessian matrix of the function $f$ evaluated in $x^{\star}$, where the element $(i,j)$ refers to $\partial_{x_{i},x_{j}}f(x^{\star}),\,\forall i,j\in\{1,\ldots,d\}$.
 TABLE I: List of variables, index and sets.
@@
 $z(n)$ Zone of producer $n$
-$\mathcal{N}(z)$ Subset of producers in zone $z$
+$\mathcal{N}(z)$ Subset of producers in $z$
 $K_{n}$ Maximum number of bids for producer $n$
@@
 II An Adversarial Game Formulation
-The ancillary services market can be formulated as a multi-leader single follower bilevel problem.
+The ancillary services market can be formulated as a multi-leader single follower Stackelberg game.
 At the upper level, producers (leaders) bid strategically to maximize their revenues while anticipating the market clearing.
 At the lower level, the market operator (follower) determines the fractions of the bids that are accepted in order to minimize the total cost it has to pay to the producers to meet the zonal demands, while satisfying zonal, interzone and bid-related constraints.
-Because of the opposite objectives of the producers and market operator, the problem can be interpreted as an adversarial game, also called a zero sum game.
+Because of the opposite objectives of the producers and market operator, the problem can be interpreted as an adversarial game, also called a zero-sum game.
 We define $\mathcal{N}:=\{1,\ldots,N\}$ as the set of $N$ producers and $\mathcal{Z}:=\{1,...,Z\}$ as the set of zones.
@@
 For all $n\in\mathcal{N}$, $z(n)\in\mathcal{Z}$ defines the zone of producer $n$ and $\mathcal{N}(z)\subseteq\mathcal{N}$ is the subset of producers in zone $z$.
-Each producer $n$’s decision variable is a finite sequence of bids $y_{n}:=\text{col}(\Delta_{n},\pi_{n})=\text{col}\Big{(}(\Delta_{n,1},\pi_{n,1}),\dots,(\Delta_{n,K_{n}},\pi_{n,K_{n}})\Big{)}$, where $\Delta_{n},\pi_{n}$ are producer $n$’s finite sequences of offered capacities and prices, respectively.
+Each producer $n$’s decision variable is a finite sequence of bids
+$y_{n}:=\text{col}(\Delta_{n},\pi_{n})=\text{col}\Big((\Delta_{n,l},\pi_{n,l})_{l=1,\ldots,K_{n}}\Big),$ (1)
+where $\Delta_{n},\pi_{n}$ are producer $n$’s finite sequences of offered capacities and prices, respectively. We denote by $y:=(y_{n})_{n}$ the producers’ strategy profile.
 The decision variables of the market operator are $x:=(x_{n,k}^{z})_{n,k,z}\in[0,1]^{B\times Z}$, where $x_{n,k}^{z}$ represents the fraction of bid activated in each zone.
-A summary of the definitions is in Tab.
+A summary of the definitions is displayed in Tab.
 I.
@@
 […]
-$J_{n}(y_{n},x_{n}):=\sum_{k=1}^{K_{n}}\pi_{n,k}\Delta_{n,k}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\right).$ (2)
+$J_{n}(x_{n},y_{n}):=\sum_{k=1}^{K_{n}}\pi_{n,k}\Delta_{n,k}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\right).$ (3)
 II-B Market Operator
-The market operator deals with four sets of constraints that depend on the capacities $\Delta_{n,k},\forall k\in[1,K_{n}],n\in\mathcal{N}$ submitted by the producers:
+The market operator deals with four sets of constraints that depend on the capacities $\Delta_{n,k},\forall k\in[1,K_{n}],n\in\mathcal{N}$ submitted by the producers
 Zonal Demand
@@
 (5) captures the fact that for a given bid $(n,k)$, the sum of the fractions activated in each zone is normalized.
-Constraints (3)-(6) can be compactly written under the form
+To highlight the mathematical structure of the problem, constraints (4)-(7) can be compactly written under the form
 $\Phi(x,y)\leq 0,$
-and we denote $\mathcal{X}(y):=\{x\in[0,1]^{B\times Z}:\Phi(x,y)\leq 0\}$ the feasibility set of the market operator which depends on the profile of strategies of the producers $y:=(y_{n})_{n}$.
+where $y:=(y_{n})_{n}$, $y_{n}$ is defined in Eq.
+(1), and $\Phi:[0,1]^{B\times Z}\times\prod_{n\in\mathcal{N}}[[\underline{\Delta}_{n},\bar{\Delta}_{n}]^{K_{n}}\times[\underline{\pi}_{n},\bar{\pi}_{n}]^{K_{n}}]\to\mathbb{R}^{3Z+N}$.
+We denote
+$\mathcal{X}(y):=\{x\in[0,1]^{B\times Z}\,|\,\Phi(x,y)\leq 0\},$
+the feasible set of the market operator which depends on the profile of strategies of all the producers $y:=(y_{n})_{n}$.
 The market operator aims to minimize the total cost it has to pay to the producers.
-Its objective function takes the form:
+Its objective function takes the form
 $J_{\text{MO}}(x,y):=\sum_{n\in\mathcal{N}}\sum_{1\leq k\leq K_{n}}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\right)\Delta_{n,k}\pi_{n,k}.$ (8)
-The market operator determines the fractions of bids to activate in each zone $(x_{n,k}^{z})_{n,k,z}$, under constraints (3)-(6):
+The market operator determines the fractions of bids to activate in each zone $(x_{n,k}^{z})_{n,k,z}$, under constraints (4)-(7)
 $\min_{x}$ $\quad J_{\text{MO}}(x,y),$ (9a)
 s.t. $\quad x\in\mathcal{X}(y).$ (9b)
-We report a schematic view of a two-zone market in Fig. 1, where we highlight with arrows the different zonal constraints.
+We report a schematic view of a two-zone market in Fig. 1, where we highlight with arrows the different constraints.
+Zone AZone BZonal Demand AZonal Demand BExportCore PortionCore PortionP1P2P3P4P5
 Figure 1: Representation of an ancillary market with two zones, with three producers in the first zone and two producers in the second zone. Constraints are highlighted with arrows.
 II-C Market Clearing and KKT Reformulation
-We formulate the ancillary market clearing problem as a pessimistic multi-leader single follower bilevel optimization problem:
-$\forall n\in\mathcal{N},\quad$ $\min_{x_{n}}\max_{y_{n}}$ $\quad J_{n}(y_{n},x_{n}),$ (9a)
+We formulate the ancillary services market clearing problem as a pessimistic multi-leader single follower Stackelberg game
+$\forall n\in\mathcal{N},\quad$ $\min_{x_{n}}\max_{y_{n}}$ $\quad J_{n}(x_{n},y_{n}),$ (10a)
 s.t. $x\in\operatorname*{argmin}_{x\in\mathcal{X}(y)}\quad J_{\text{MO}}(x,y),$ (10b)
@@
 In case of multiple solutions at the lower level, the pessimistic formulation enables the selection of the worst equilibrium for the producers, obtaining a more robust formulation of the market.
-The Lagrangian function associated with the market clearing defined in problem (8) takes the form:
+The Lagrangian function associated with the market clearing defined in problem (9) takes the form
 $L(x,y;\Lambda)=\sum_{n\in\mathcal{N}}\sum_{k=1}^{K_{n}}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\right)\Delta_{n,k}\pi_{n,k}+\Lambda^{\top}\Phi(x,y),$ (11)
-where $\Lambda\in\mathbb{R}_{+}^{3Z+B}$ refers to the dual variables associated with constraints (3)-(6). The analytical expression of the Lagrangian function in (10) can be found in Appendix A.
+where $\Lambda\in\mathbb{R}_{+}^{3Z+B}$ refers to the dual variables associated with constraints (4)-(7).
 Lemma 1 (Slater’s Constraint Qualification).
@@
     ${\sum_{z^{\prime}\neq z}}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\bar{\Delta}_{n}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}$
-    and $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$.
+    and
+    $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$.
 Proof.
 Under condition (i), each zone can satisfy its own demand; thus, there exists an interior-point solution.
-Under condition (ii), one zone $z\in\mathcal{Z}$ may not be able to meet its own demand, i.e., $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}>0$.
-In this situation, we need enough demand offers left from other zones to cover the lack of supply, thus $\sum_{z^{\prime}\neq z}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\bar{\Delta}_{n}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}$.
+Under condition (ii), one zone $z\in\mathcal{Z}$ may not be able to meet its own demand, i.e.
+$D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}>0.$
+In this situation, we need enough demand offers left from other zones to cover the lack of supply, thus
+$\sum_{z^{\prime}\neq z}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\bar{\Delta}_{n}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}.$
 Since the export constraint should hold, we get that $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}.$
-∎
 When the lower-level problem (9) satisfies Lemma 1, we can replace it by its KKTs in (10).
-Hence, each producer has to anticipate the decision $x$ of the market operator, meaning that in the single level the decision variables for producer $n$ will be $y_{n}\in\mathcal{Y}_{n}$ and $x_{n}\in\mathcal{X}_{n}(y)$.
-Before writing the full reformulation, we introduce tracking variables for every $z\in\mathcal{Z}$:
+Hence, each producer has to anticipate the decision $x$ of the market operator, meaning that in the single level the decision variables for producer $n$ will be $y_{n}\in\mathcal{Y}_{n}$ and $x_{n}\in\mathcal{X}_{n}(y)$ with $y=(y_{n})_{n}$.
+Before writing the full reformulation, we introduce tracking variables for every zone $z\in\mathcal{Z}$
 $d^{z}_{n}(x_{-n},y_{-n}):=D_{z}-\sum_{j\neq n}\sum_{1\leq k\leq K_{j}}x_{j,k}^{z}\Delta_{j,k},$ (12a)
@@
 $c^{z}_{n}(x_{-n},y_{-n}):=C_{z}-\sum_{\begin{subarray}{c}j\in\mathcal{N}(z),\\ j\neq n\end{subarray}}\sum_{1\leq k\leq K_{j}}x_{j,k}^{z}\Delta_{j,k}.$ (12c)
-The single-level problem for agent $n$ is
+The single-level problem for agent $n$ takes the form
 $\min_{x_{n}}\,\max_{y_{n},\Lambda_{n}}J_{n}(y_{n},x_{n}),$ (13a)
 $\mbox{s.t. }y_{n}\in\mathcal{Y}_{n},x_{n}\in\mathcal{X}_{n}(y),\Lambda_{n}\geq 0,$ (13b)
-$\nabla_{x_{n}}L(x_{n},y_{n};\Lambda_{n})=0,$ (12c)
+$\nabla_{x_{n}}L(x,y;\Lambda)=0,$ (13c)
 $\lambda_{n}^{z}\Bigg(\sum_{1\leq k\leq K_{n}}x_{n,k}^{z}\Delta_{n,k}-d_{n}^{z}(x_{-n},y_{-n})\Bigg)=0,\forall z\in\mathcal{Z},$ (13d)
@@
 (13d), (13f), (13g) impose couplings through bids and zones.
-We now introduce a more compact notation.
-First we stress that, given the dependence of the feasibility set $\mathcal{X}$ on $y$, the decision variable $x_{n}$ can be expressed as an implicit function in $y$, leading to $x_{n}\equiv x_{n}(y)$.
-This in turn, allows us to write the objective function as an explicit function of all producers’ decision variables, i.e.
-$J_{n}(y_{n},y_{-n})$.
-Then we let $w_{n}:=(\pi_{n,k},\Delta_{n,k},x_{n,k}^{z},\lambda_{n}^{z},\mu_{n,k},\sigma_{n},\delta_{n})\in\mathcal{W}_{n}:=[\underline{\pi}_{n},\bar{\pi}_{n}]^{K_{n}}\times[\tilde{\Delta}_{n},\bar{\Delta}_{n}]^{K_{n}}\times[0,1]^{K_{n}\cdot Z}\times[0,+\infty]^{Z}\times[0,+\infty]^{K_{n}}\times[0,+\infty]\times[0,+\infty]$.
-$\mathcal{W}_{n}$ is a polyhedron with dimensionality $3K_{n}+Z(K_{n}+1)+2$.
-The set of $l_{n}$ local constraints are captured by $h_{n}:\omega_{n}\rightarrow\mathbb{R}^{l_{n}}$ and that of the $r$ shared constraints by $G:\Omega\rightarrow\mathbb{R}^{r}$, where $\Omega_{n}\supseteq\mathcal{W}_{n}$ and $\Omega:=\prod_{n=1}^{N}\Omega_{n}$.
-In particular $h_{n}$ and $G$ are continuously differentiable mappings.
-Finally, we introduce the local feasibility set for each player $n$: $\widetilde{\mathcal{W}}_{n}:=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0\}$.
-Using the previously introduced compact notation with the explicit objective function, we get $J_{n}(w_{n},w_{-n})$.
-Thus, the $N$ optimization problems in Eqs.
-(12) give rise to a generalized Nash game $\Gamma$ that we denote in strategic form $\Gamma:=(\mathcal{N},\prod_{n}\mathcal{W}_{n},(J_{n})_{n})$:
+We now aim to highlight the structure of the problem.
+We first note that, given the dependence of the feasible set $\mathcal{X}$ on $y$, the decision variable $x_{n}$ can be expressed as an implicit function in $y$, $x_{n}(y)$.
+This allows to write the producer’s objective function as an explicit function of all producers’ decision variables, i.e., $J_{n}(y_{n},y_{-n})$.
+However, in this reformulation, the dual variables are decision variables of the producers. To simplify, we need an additional change of variable. We set
+$w_{n}:=(\pi_{n,k},\Delta_{n,k},x_{n,k}^{z},\lambda_{n}^{z},\mu_{n,k},\sigma_{n},\delta_{n})\in\mathcal{W}_{n},$
+where $\mathcal{W}_{n}:=[\underline{\pi}_{n},\bar{\pi}_{n}]^{K_{n}}\times[\underline{\Delta}_{n},\bar{\Delta}_{n}]^{K_{n}}\times[0,1]^{K_{n}\cdot Z}\times[0,+\infty]^{Z}\times[0,+\infty]^{K_{n}}\times[0,+\infty]\times[0,+\infty]$ is a polyhedron of dimension $3K_{n}+Z(K_{n}+1)+2$.
+The set of the $l_{n}$ local constraints are represented by the function $h_{n}:\mathcal{W}_{n}\rightarrow\mathbb{R}^{l_{n}}$. Further, we introduce the local feasible set for each player $n$
+$\widetilde{\mathcal{W}}_{n}:=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0\}.$
+Similarly, the $r$ shared constraints are formulated in a compact form as $G:\Omega(w)\rightarrow\mathbb{R}^{r}$, where $\Omega_{n}(w_{-n}):=\{w_{n}\in\widetilde{\mathcal{W}}_{n}\,|\,G(w_{n},w_{-n})\leq 0\}$ and $\Omega(w):=\prod_{n=1}^{N}\Omega_{n}(w_{-n})$.
+We check that $h_{n}$ and $G$ are continuously differentiable mappings.
+Relying on these notations, the $N$-producer coupled optimization problems, derived from Eqs. (13), give rise to a nonconvex game with side constraints [pang_nonconvex_2011]
+$\Gamma:=\Big(\mathcal{N},\Omega(\cdot),\left(J_{n}(\cdot)\right)_{n}\Big),$
+where producers solve the parametrized optimization problems
 $\forall n\in\mathcal{N},\quad$ $\max_{w_{n}\in\mathcal{W}_{n}}$ $J_{n}(w_{n},w_{-n}),$ (14a)
@@
 $G(w_{n},w_{-n})\leq 0.$ (14c)
+Pang and Scutari in [pang_nonconvex_2011] study the properties of problem (14), providing different sets of assumptions to be used in order to obtain existence guarantees for specific equilibria.
+In the next section, we rely on a result guaranteeing the existence of a Nash equilibrium for a relaxed version of the problem.
+Further, under mild assumptions on the model parameters, we show the equivalence between the set of solutions of the relaxed version and the generalized Nash game $\Gamma$.
+Refinements of generalized Nash equilibria (GNEs) solutions of $\Gamma$ requiring the uniform evaluation of the Lagrange multipliers associated with the coupling constraints across the producers, are called variational equilibria (v-GNEs) [harker1991].
+Interestingly, the generalized potential structure of $\Gamma$, guarantees the economic efficiency of v-GNEs.
+However, reaching v-GNEs requires some implicit coordination between the market participants which may be difficult to enforce.
+This will be discussed further in Section V.
 III Theoretical Framework
 In a large part of the literature, the convexity of the agents’ optimization problems is an essential assumption under which noncooperative games are studied and analyzed.
-On the contrary, in this paper, $\Gamma$ is a generalized Nash game with side constraints, involving trilinear objective functions and nonconvex constraints thus giving rise to a nonconvex game.
-In order to proceed we will rely on results established in [18], where the authors study a class of problems with similar settings.
-To that purpose, we first introduce a penalized version of our problem, which is more tractable for this analysis, providing results regarding the equivalence with the non-penalized version; then we present the settings of [18], which we use to prove existence of market equilibrium under mild assumptions.
-Finally we highlight the potential structure of $\Gamma$.
-III-A Penalized Generalized Nash Game
-We consider a penalized version of our problem, which makes it more tractable and it will allow us to use mild assumptions to obtain an equilibrium existence guarantee.
-We add the constraints in Eqs.
-(3) and (6) to the objective function, which gives the function $J_{n}^{p}(w)$ and the new constraint maps $h_{n}^{p}(w_{n})$ and $G^{p}(w)$ (see Appendix B for more details).
+On the contrary, in this paper, $\Gamma$ is a generalized Nash game with side constraints, involving trilinear objective functions and nonconvex constraints; thus, giving rise to a nonconvex game.
+In order to proceed, we will rely on results established in [pang_nonconvex_2011], where the authors study a class of problems with similar settings.
+To that purpose, we first apply the generic setting of [pang_nonconvex_2011].
+Then, we present a relaxed version of our problem, which is more tractable for this analysis, providing results regarding the equivalence with the non-penalized version.
+Further, we prove existence of market equilibrium under mild assumptions, and highlight the generalized potential structure of $\Gamma$.
+III-A Nonconvex Game with Side Constraints
+The formulation of [pang_nonconvex_2011] considers a problem where the coupling constraints are treated as penalties with associated prices, in a modified objective function.
+This problem is then solved as an $(N+1)$-player game, where $N$ players optimize their variables and an external one optimizes the prices, which are shared by all players.
+In our formulation, the coupling constraints are Eqs.
+(4)-(6)-(7)-(13d)-(13f)-(13g), while the remaining part is considered as local constraints.
+With the compact notation previously introduced, we can write the problem as an $(N+1)$-player game with prices
+$\forall n\in\mathcal{N},\quad\max_{w_{n}\in\widetilde{\mathcal{W}}_{n}}\left[J_{n}(w_{n},w_{-n})-\sum_{j=1}^{r}\rho_{j}G_{j}(w)\right],$ (15)
+where we call the externally optimized price, $\rho\in\mathbb{R}^{r}_{+}$. We can now recall the definition of Nash Equilibrium.
+Definition 1 (Nash Equilibrium with Side Constraints, [pang_nonconvex_2011, Definition 1]).
+A Nash equilibrium (NE) for the game (15) is a strategy-price tuple $(w^{\star},\rho^{\star})$, such that for all $n$
+$w_{n}^{\star}\in\operatorname*{argmax}_{w_{n}\in\widetilde{\mathcal{W}}_{n}}\left[J_{n}(w_{n},w_{-n}^{\star})-\sum_{j=1}^{r}\rho_{j}^{\star}G_{j}(w_{n},w_{-n}^{\star})\right],$
+and
+$0\leq\rho^{\star}\perp G(w^{\star})\leq 0.$
+III-B Relaxed Generalized Nash Game
+In order to extend the results from [pang_nonconvex_2011], we consider a relaxed version of our problem.
+We will use this relaxed version to prove the existence of an equilibrium under mild assumptions.
+Concretely, we add the constraints from Eqs.
+(4) and (7) to the objective function, which gives the function $J_{n}^{p}(w)$, the new constraint maps $h_{n}^{p}(w_{n})$ and $G^{p}(w)$, and the associated set of penalties $M_{n}$ for each agent.
 We can formulate the penalized generalized Nash game $\Gamma^{p}$ as follows:
-$\forall n\in\mathcal{N},\quad$ $\min_{w_{n}\in\mathcal{W}_{n}}$ $-J_{n}^{p}(w)$ (14a)
+$\max_{w_{n}\in\mathcal{W}_{n}}$ $J_{n}^{p}(w)$ (16a)
 s.t. $h_{n}^{p}(w_{n})\leq 0$ (16b)
@@
 Proposition 1.
-If $K_{n}\geq\frac{Z}{3}$, there exists a set of penalties such that a solution of the penalized game (14) is a solution of $\Gamma$.
-Proposition 1 results from the comparisons of the sets of KKTs of the generalized Nash game $\Gamma$ and the penalized game (14) denoted $\Gamma^{p}$. Details can be found in Appendix C.
-Thanks to this reformulation we obtain copositive Hessians matrices associated to the constraints $h_{n}^{p}(w_{n})$ and $G^{p}(w)$, which is the key point in obtaining an existence result.
-We discuss the form of these matrices in Appendix D.
-III-B Nonconvex Game with Side Constraints
-The formulation of [18] considers a problem where the coupling constraints are treated as penalties with associated prices, in a modified objective function.
-This problem is then solved as an $(N+1)$-player game, where $N$ players optimize their variables and an external one optimizes the prices, which are shared by all players.
-With the compact notation previously introduced, we can write the problem in Eq.
-(14) as an $(N+1)$-player game with prices [^1] Notice that to directly extend the results of [18], we now work with the negative of the objective function and in minimization settings.:
-$\forall n\in\mathcal{N},\quad\min_{w_{n}\in\widetilde{\mathcal{W}}^{p}_{n}}\left[-J_{n}^{p}(w_{n},w_{-n})+\sum_{j=1}^{r}\rho_{j}G_{j}^{p}(w)\right]$ (15)
-where we call the externally optimized price, $\rho\in\mathbb{R}^{r}_{+}$ and introduce the penalized local feasibility set $\widetilde{\mathcal{W}}_{n}^{p}:=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}^{p}(w_{n})\leq 0\}$.
-We can now recall the definition of Nash Equilibrium [18, Definition 1] that is a relevant solution concept to analyze this class of games:
-Definition 1 (Nash Equilibrium with Side Constraints).
-A Nash equilibrium (NE) for the game (15) is a strategy-price tuple $(w^{\star},\rho^{\star})$, such that for all $n$:
-$w_{n}^{\star}\in\operatorname*{argmin}_{w_{n}\in\widetilde{\mathcal{W}}_{n}}\left[-J_{n}(w_{n},w_{-n}^{\star})+\sum_{j=1}^{r}\rho_{j}^{\star}G_{j}(w_{n},w_{-n}^{\star})\right],$
-and
-$0\leq\rho^{\star}\perp G(w^{\star})\leq 0.$
+Assume $K_{n}\geq\frac{Z}{3},\,\forall n$. Then, there exists a set of penalties $(M_{j})_{j\in\mathcal{N}}$ such that a solution of the penalized game (16) is a solution of $\Gamma$.
+Proof.
+We call the two Lagrangian functions associated to the problems $\mathcal{L}_{n}^{P}$ and $\mathcal{L}_{n}^{U}$, and we consider the solution for $U$: $(w_{n}^{*},\Theta_{n}^{*})\in\mbox{SOL}\{\mbox{KKT of }U\}$, where $\Theta$ is the set of Lagrangian multipliers.
+This solution must respect all the constraints and be a solution of $\nabla\mathcal{L}^{U}(w^{*}_{n},\Theta^{*}_{n})=0$.
+We now call the part of Lagrangian multipliers that are still present in $P$ $\Theta[\Sigma_{n}]$, and we consider the set of equations coming from $\nabla\mathcal{L}^{P}(w_{n},\Theta_{n}[\Sigma_{n}];M_{n})=0$, where $M_{n}$ is the set of penalties for agent $n$.
+If we evaluate these equations in $(w_{n}^{*},\Theta_{n}^{*}[\Sigma_{n}])$, we can solve them w.r.t.
+the penalties $M_{n}$, obtaining $\bar{M}_{n}$.
+Given the assumptions, solving the system $\nabla\mathcal{L}^{P}(w_{n},\Theta_{n}[\Sigma];\bar{M}_{n})=0$ in $(w_{n},\Theta_{n}[\Sigma])$ provides $\tilde{w}_{n}$ which is a solution also for $\nabla\mathcal{L}^{U}(w_{n},\Theta_{n})=0$.
+∎
 III-C Existence of a Generalized Nash Equilibrium for $\Gamma$
-To obtain existence result for the penlaized problem, we have to verify assumptions on the variables’ domain, the objective functions and the constraint sets.
-Thanks to our penalized reformulation, we only need to verify a set of mild assumptions:
+To obtain existence result for the relaxed problem, we have to verify assumptions on the variable domains, the objective functions and the constraint sets.
 Assumptions 1.
@@
   • (b)
-    $J_{n}(\cdot,w_{-n})$ is continuously differentiable in $w_{n}$, $h_{n}(\cdot)$ and $G(\cdot,w_{-n})$ are twice continuously differentiable in $w_{n}$ for all $n$;
+    $J_{n}^{p}(\cdot,w_{-n})$ is continuously differentiable in $w_{n}$, $h_{n}^{p}(\cdot)$ and $G^{p}(\cdot,w_{-n})$ are twice continuously differentiable in $w_{n}$ for all $n$;
   • (c)
-    There exists $w^{\sharp}\in\mathcal{W}:=\prod_{n\in\mathcal{N}}\mathcal{W}_{n}$ such that $h_{n}(w_{n}^{\sharp})<0,\,\forall n\in\mathcal{N}$ and $G(w^{\sharp})<0$;
-Theorem 1 (Existence of a solution for $\Gamma$).
+    There exists $w^{\star}\in\mathcal{W}:=\prod_{n\in\mathcal{N}}\mathcal{W}_{n}$ such that $h_{n}^{p}(w_{n}^{\star})<0,\,\forall n$ and $G^{p}(w^{\star})<0$;
+  • (d)
+    The Hessians matrices $\nabla^{2}h_{n}^{p}(w_{n})$ for the $l_{n}$ local constraints are copositive for all $n$;
+  • (e)
+    The Hessians matrices $\nabla^{2}G^{p}(w)$ for the $r$ global constraints are copositive.
+Theorem 1 (Existence of a solution for $\Gamma^{p}$).
 Define $\mathcal{W}^{\star}\subseteq\mathcal{W}$ such that for any $w\in\mathcal{W}^{\star}$, the Hessian matrix of the objective function is positive semi-definite.
-Then, there exists a generalized Nash equilibrium on $\mathcal{W}^{\star}$ solution of $\Gamma^{p}$.
+Then, there exists a generalized Nash equilibrium on $\mathcal{W}^{\star}$ solution of $\Gamma$.
 Proof.
@@
 Assumption 1 (c) requires the existence of an interior point.
-Considering $w^{\sharp}$ that strictly satisfies Eq.
-(3), (4), (5) and (6), we immediately obtain that $\lambda_{n}^{z}=0,\mu_{n,k}=0,\sigma_{n}=0,\delta_{n}=0,\,\forall k\in[1,K_{n}],\,\forall n\in\mathcal{N},\,\forall z\in\mathcal{Z}$.
+Considering $w^{\star}$ that strictly satisfies Eq.
+(4), (5), (6) and (7), we immediately obtain that $\lambda_{n}^{z}=0,\mu_{n,k}=0,\sigma_{n}=0,\delta_{n}=0,\,\forall k\in\{1,\ldots,K_{n}\},\,\forall n\in\mathcal{N},\,\forall z\in\mathcal{Z}$.
 Eq.
 (13c) needs instead to be relaxed to a strict inequality.
-We then apply [18, Theorem 13] to the $(N+1)$-players game to obtain the existence of a generalized Nash equilibrium for $\Gamma^{p}$.
-∎
-Corollary 1.
-The generalized Nash game $\Gamma$ defined in (13) admits a solution.
-Proof.
-Using results from [18, Sec.
-2.1], we have an equivalence between the Nash equilibrium of the $(N+1)$-players game and the associated generalized Nash game, which in our case is the penalized version $\Gamma^{p}$.
-We conclude using Proposition 1, which states the equivalence between $\Gamma^{p}$ and $\Gamma$.
+Finally, to check for Assumptions 1 (d)-(e), we need to compute the Hessians matrices associated to the constraints functions.
+Regarding the local set of constraints $h_{n}^{p}(w_{n})$, we obtain $\nabla^{2}h_{n}^{p}(w_{n})=0,\,\forall n\in\mathcal{N},\,\forall w_{n}\in\mathcal{W}_{n}$, except for Eq.
+(13c), for which we obtain the following cases:
+$x_{n}^{\top}\nabla_{x_{n}}^{2}\left(\nabla_{x_{n,k}^{z}}L\right)x_{n}=\begin{cases}\pi_{n,k}-\lambda_{n}^{z}-\delta_{n}\geq 0,\quad\mbox{if }z=z(n)\\ \pi_{n,k}-\lambda_{n}^{z}+\sigma_{n}\geq 0,\quad\mbox{if }z\neq z(n)\end{cases}$
+Given that an interior point forces $\lambda_{n}^{z}=0,\delta_{n}=0$ and $\sigma_{n}=0$, and that the prices $\pi_{n,k}$ are positive by definition, both inequalities are always satisfied.
+Considering the only remaining global constraint, Eq (6), we obtain a constant positive Hessian , which directly satisfies the copositivity assumption.
+We then apply [pang_nonconvex_2011, Theorem 13] to the $(N+1)$-players game to obtain the existence of a Nash equilibrium $(w^{\star},\rho^{\star})$, and from [pang_nonconvex_2011, Section 2.1] we can conclude that $w^{\star}$ is a generalized Nash equilibrium for $\Gamma^{p}$.
+Finally, using Proposition 1, a solution of $\Gamma^{p}$ is also a solution for $\Gamma$, which concludes the proof.
 ∎
@@
 Proof.
-Using [19, Def.
+Using [facchinei_decomposition_2011, Def.
 2.1], two requirements need to be satisfied for a GNG to be a GPG.
-The first requirement amounts to have a feasible set that is nonempty, which is easily satisfied in our setting.
-In particular, the feasible set, which we call $C$, does not have to be the cartesian product of the players’ feasible sets, meaning that $C\subseteq\prod_{n=1}^{N}\bar{\mathcal{W}}_{n}(w_{-n})$, where $\bar{\mathcal{W}}_{n}(w_{-n}):=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0,G(w)\leq 0\}$.
-The second requirement is for a potential $P(w)$ function to exist, which in our case is satisfied by using $P(w):=\sum_{n=1}^{N}J_{n}(w)$.
+The first requirement amounts to have a feasible set that is nonempty, which in our setting implies having enough energy offer to cover the total demand.
+The second requirement is for a potential $P(w)$ function to exist, which in our case is satisfied by using $P(w):=\sum_{n=1}^{N}J_{n}(w_{n},w_{-n})$.
 ∎
-From Proposition 2, we can rewrite Eqs. (13) as:
+From Proposition 2, we can rewrite Eqs. (14) as
 $\max_{w\in W}$ $\quad\sum_{m=1}^{N}J_{n}(w_{m},w_{-m}),$ (17a)
 s.t. $\quad h_{n}(w_{n})\leq 0,\forall n,$ (17b)
-$\quad G(w_{n},w_{-n})\leq 0,\,\forall n,$ (16c)
-which can be solved using a standard nonlinear programming solver.
+$\quad G(w_{n},w_{-n})\leq 0,$ (17c)
+which can be solved using a standard nonlinear programming solver. The potential structure is interesting algorithmically, to guarantee the convergence of best-response dynamics.
 IV Algorithmic Solutions
-We solve the market problem relying on integrated optimization and Gauss-Seidel best-response method, that we compare against a MARL algorithm.
-For the integrated optimization approach we use the MINOS solver, which relies on a projected Lagrangian method [20].
-Next, we discuss the best-response algorithm we use and how we built the multi-agent simulation.
-IV-A Gauss-Seidel Best Response
-We follow the implementation of the Gauss-Seidel Best Response algorithm from [19].
-Here we remove the coordination between players, and we solve the problem in Eq.
-(13) for each agent iteratively (once again with the MINOS solver), stopping the whole process once the decision variables of the agents do not change anymore, as summarized in Alg.
+We propose two exact approaches to compute GNEs of $\Gamma$, that we compare against MARL simulation which capture in a realistic way the interactions between the producers and the market dynamics.
+One of the exact approach makes use of the GPG structure of $\Gamma$; thus leading to compute a solution of (17) in an integrated way, relying on a nonlinear mathematical optimization solver (MINOS) which is based on the projected Lagrangian method [noauthor_minos_nodate].
+We describe next the other exact approach.
+IV-A Gauss-Seidel Best-Response
+We follow the implementation of the Gauss-Seidel best-response algorithm from [facchinei_decomposition_2011].
+We solve the problem in Eq.
+(14) for each agent sequentially (using MINOS to compute best-response for each player); stopping the whole process once the stopping criterion is met, i.e., the distance between the players’ strategy profiles at two consecutive iterations do not change more than a small constant $\varepsilon$.
+This is summarized in Alg.
 1.
 Algorithm 1 Gauss-Seidel Best Response.
-1:Initialize problem with feasible point $w^{0}=[w_{1}^{0},\dots,w_{N}^{0}]$, $k=0$
-2:while Termination criterion is false do
-3:  for $n\in[1,\dots,N]$ do
+1:Initialize problem with feasible point $w^{0}=[w_{1}^{0},\dots,w_{N}^{0}]$, $k=0$, $\varepsilon>0$
+2:while $w^{k}-w^{k-1}||>\varepsilon$ do
+3:  for $n\in\{1,\dots,N\}$ do
 4:   Solve regularized Eq. (14) for agent $n$ with $w_{1}^{k+1},\dots,\allowbreak w_{n}^{k},\dots,\allowbreak w_{N}^{k+1}]$
@@
 8:end while
-The integrated solution can be seen as a coordinated solution for all the producers at once, while using the Best-Response algorithm allows for each producer to optimize its strategy alone.
-In the results section we will explain more in details how these two solutions differ in modeling the interactions between agents.
-Furthermore, from [19, Theorem 5.2] and the potential structure of our game, we have a guarantee of convergence to a Nash equilibrium, provided it exists, which we already proved in Section III-C.
-These two approaches allow us to compare fully coordinated and uncoordinated interaction schemes between agents, that we will compare against the multi-agent simulation based on reinforcement learning.
-IV-B Multi Agent Deep Reinforcement Learning
+From [facchinei_decomposition_2011, Theorem 5.2] and the GPG structure of the game, the Gauss-Seidel best-response algorithm is proved to converge in finite time to a GNE of $\Gamma$.
+IV-B Multi-Agent Deep Reinforcement Learning
 The core idea of reinforcement learning is to have an agent learning directly from exposure to an unknown environment.
@@
 We define the discounted return over a trajectory of length $L$ as $R(s,a):=\sum_{t=0}^{t=L-1}\gamma^{t}r(s_{t},a_{t})$, where $0<\gamma<1$ is the discount factor.
-An agent interacts with the environment sampling action from its policy $a\sim\mu(s)$, and it aims at maximizing the reward, so the goal is to find the policy that maximizes the discounted return over an infinite horizon.
+An agent interacts with the environment sampling action from its policy $a\sim\mu(s)$, and it aims to maximize the reward, so the goal is to find the policy that maximizes the discounted return over an infinite horizon.
 In particular, defining the $Q$-function as:
 $Q_{\mu}(s,a)=\mathbb{E}[R(s,a)|\mu,s_{0}=s,a_{0}=a],$
-we want to solve the following optimization to obtain the optimal policy $\pi^{*}$:
-$Q^{*}(s,a)=Q_{\mu^{*}}(s,a)=\max_{\mu}Q_{\mu}(s,a).$
-Generally, we do not have access to analytical form for neither $Q$ nor $\mu$, making the optimization extremely complex to solve.
+we want to solve the following optimization problem to obtain the optimal policy $\pi^{\star}$
+$Q^{\star}(s,a)=Q_{\mu^{\star}}(s,a)=\max_{\mu}Q_{\mu}(s,a).$
+Generally, we do not have access to analytical forms for neither $Q$ nor $\mu$, making the optimization extremely complex to solve.
 RL algorithms, and in particular deep RL (DRL) algorithms, try to solve this problem learning a policy that approximates the optimal one.
@@
 We will not explore in depth here the differences between DRL algorithms, but we will provide a quick overview of the one we use.
-We implemented the multi-agent deep reinforcement learning algorithm described in [21]: this is called an actor-critic algorithm, where an actor learns the policy and a critic the $Q$-function.
+We implemented the multi-agent deep reinforcement learning algorithm described in [lowe_multi-agent_2017]: this is called an actor-critic algorithm, where an actor learns the policy and a critic the $Q$-function.
 The actor and critic are neural networks parameterized by $\theta$ and $\omega$ respectively, which are trained over specific losses.
 In particular the loss function for the critic is based on the Bellman’s equation:
-$\mathcal{L}_{Q}=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[(Q_{\omega}(s,a)-(r+\beta Q_{\omega}(s^{\prime},\mu_{\theta}(s^{\prime}))))^{2}\right],$
+$\mathcal{L}_{Q}=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[(Q_{\omega}(s,a)-(r+\gamma Q_{\omega}(s^{\prime},\mu_{\theta}(s^{\prime}))))^{2}\right],$
 where $\mathcal{D}$ is a buffer where we save previously encountered states, the actions that the agent used and the reward it obtained. The actor instead minimizes the following loss:
@@
 The intuition behind this algorithm is that the critic learns to map future rewards to specific states and actions (which is the $Q$-function), while the actor tries to maximize the $Q$-function itself (represented by the critic output).
-In the multi-agent settings we appply a modified version of the single agent implementation presented here, where each agent has its own actor and critic, but the critics receive as input the actions of all the agents.
-This was introduced in [21] as a way of stabilizing the learning and improving the results.
+In the multi-agent settings we apply a modified version of the single agent implementation presented here, where each agent has its own actor and critic, but the critics receive as input the actions of all the agents.
+This was introduced in [lowe_multi-agent_2017] as a way of stabilizing the learning and improving the results.
 In particular, given the zonal nature of our environment, we share the actions between the agents of the same zone, but not between all agents.
+In our environment, the action $a$ of each producer corresponds to its bids (prices and capacities), while the state $s$ contains the previous day clearing prices, information about the current day and a random zonal perturbation.
+The reward is a function of the profit and clearing price.
+In order to run the simulation, each producer outputs their bids, which are used to clear the market solving problem (9).
+The results are then used to compute the reward for each producer, which in turn is used to update their strategy.
 In Alg.
 2 we report in pseudo-code the structure for the training of the RL agents.
-We leave further details on the implementation and references to the code in Appendix E.
 Algorithm 2 Basic structure of the zonal MARL algorithm.
@@
 We now discuss the numerical results of the three algorithmic solutions we propose.
-We test these algorithms on real data from the German and Austrian ancillary service markets, for the period between January 2024 and August 2024 [22].
+We test these algorithms on real data from the German and Austrian ancillary services markets, for the period between January 2024 and August 2024 [noauthor_accueil_nodate].
 The maximum length for the bid sequence is 5, and each submitted capacity must be $\geq 5$ MW.
@@
 Finally, we test multiple values for the export constraints, in order to study the effect that the two zones have on each other.
-All simulations were done on a MacBook Pro with an M2 Pro chip and 16G of memory.
+All simulations were done on a MacBook Pro with an M2 Pro chip and 16G of memory (code available at https://github.com/FrancescoMorri/MARL_ancillary_markets).
 V-A Convergence of the Multi-Agent RL Simulation
@@
 Our settings are particularly complicated, since the market environment is dependent on all agents decisions.
-We tested two settings for the MARL algorithm: one with 4 learning based producers, and one with 8.
+We tested two settings for MARL algorithm: one with 4 learning based producers, and one with 8.
 In the first case, the 4 other producers always bid their marginal price.
@@
 To assess the stability and convergence, we consider the reward of each agent throughout the training.
-More precisely, we compute the average reward during an episode, this being a full pass through the data.
+More precisely, we compute the average reward during an episode, which refers to the amount of steps needed to go over the full dataset [noauthor_accueil_nodate].
 With this data we obtain Fig.
 2, where we plot the episodic average reward for each agent divided by zone, considering the case with only RL based producers.
-Notice that the reward is not the profit, as we mention in the full description of the algorithm in the Appendix.
+Notice that the reward is not the profit, as we mentioned in Section IV-B.
 From Fig.
@@
 These results are particularly interesting, since they show how the algorithm is well suited for a medium scale simulation with competitive learning agents, providing results that we can compare with the classical solutions.
-We want to stress here that, using the MARL algorithm, we compute a stationary solution for the market, without a guarantee of it being an equilibrium.
-V-B MARL Solution and Classical Solutions
-We first discuss the theoretical difference between the integrated solution and the Best-Response solution.
-Starting from the potential problem, we consider the role of the stationarity condition coming from Eq.
-(10), which for player $n$, bid $k$ and zone $z$ is the following:
-$\begin{split}&\nabla_{x_{n,k}^{z}}L(x;\lambda,\mu,\sigma,\delta)=\Delta_{n,k}\pi_{n,k}-\lambda_{z}\Delta_{n,k}+\mu_{n,k}+\\ &+(1-\mathbb{I}_{z=z(n)})\sigma_{z(n)}\Delta_{n,k}-\mathbb{I}_{z=z(n)}\delta_{z(n)}\Delta_{n,k}.\end{split}$ (17)
-Notice that every $\mu$ must be always 0, since every other term is multiplied by $\Delta$ and can be grouped.
-Furthermore, $\delta$ will be 0 for most realistic cases, since the core portion is usually a small fraction of the total demand, meaning that the constraint is always saturated.
-$\sigma$ on the other hand may be different from 0, which means that $\lambda$ must be different from 0 too, given the signs in the equation.
-Lastly, we highlight that $\pi$ is indexed by agent and bid, while $\lambda$ by zone, which means that, in most cases, the prices have to be equal to $\lambda$, virtually creating a pay-as-clear market.
-On the other hand, when using the Best Response algorithm, every agent computes its own $\lambda$, meaning that they are not forced to bid at the same price, and they can actually bid at very high prices with no repercussions.
-This constraint creates two very different markets, providing two different solutions to analyze.
-Regarding the solution obtained from the MARL algorithm, even if there are less theoretical guarantees, we can consider it as a more realistic simulation of the market, since every producer has access only to its own information and its only aim is to increase its profit.
-We start by analyzing the average cost to run the market.
-This value is computed by summing the average profits of each producer over the full data available.
-In particular, for the MARL producers, we consider the cost for the final episode of their training, where we assume they reached stable bidding strategies.
+V-B MARL Solution and Generalized Nash Equilibria
+We first discuss the theoretical difference between the integrated solution and the Gauss-Seidel best-response solution.
+The two approaches may yield different equilibria: the integrated potential formulation leads to a v-GNE, whereas the best-response produces a GNE that may not be a v-GNE [^1] Indeed, the set of v-GNEs solutions of $\Gamma$ is included in the set of GNE solutions of $\Gamma$..
+In case a v-GNE is reached, the market allocates the demand in Eq.
+(4) through a single price system, therefore leading to the determination of one price $\lambda$ for the constraint, obtaining a pay-as-clear market.
+On the contrary, in case there is no market to determine the price system associated with the coupling constraints, two producers $n$, $m$ might attribute different evaluations to the same constraint.
+This can lead to different valuations of the dual variables, in particular $\lambda_{n}\neq\lambda_{m}$, creating a virtual pay-as-bid market.
+The solution obtained via MARL simulation is stationary with respect to the reward functions used by the agents.
+There is no guarantee that this is an exact equilibrium, but the aim is to simulate the agent interactions under varying information-sharing scenarios to analyze strategic behaviors and their impacts on market dynamics.
+We first analyze the average cost to run the market, computed by summing the average profits of each producer over the full data available.
+In particular, for MARL producers, we consider the cost for the final episode of their training, where we assume they reached stable bidding strategies.
 The results are reported in Fig.
 3.
-We can immediately see how the two exact algorithms produce higher costs than the learning-based counterpart, which can be explained by the form of the constraints, as detailed above, and by the lack of actual competition.
+The two exact algorithms produce higher costs than the learning-based counterpart, which can be explained by the form of the constraints, as detailed above, and by the lack of actual competition.
 In the integrated solution, the problem is solved by an exogenous external coordinator, so by definition there will be no competition; while for the best response, each producer is free to update its decision variables uncoupled from the others.
-We can see the average cost for the MARL simulation with only 4 learning producers is very low, but this is also affected by the presence of static producers that always bid at marginal cost, driving the electricity price down.
+We can see the average cost for MARL simulation with only 4 learning producers is very low, but this is also affected by the presence of static producers that always bid at marginal cost, driving the electricity price down.
 Figure 3: Average cost of running the market, obtained with different algorithms.
 It is interesting to analyze how the profit is shared between producers in the different solutions.
-We do so by computing the Gini index [23], which, given a set of profits $p_{1}\leq p_{2}\leq\dots\leq p_{n}$, can be obtained as:
-$g=\frac{1}{n}\left(n+1-2\left(\frac{\sum_{i=1}^{n}(n+1-i)p_{i}}{\sum_{i=1}^{n}p_{i}}\right)\right).$
-The lower $g$ is, the more equal is the sharing of profits.
+To that purpose, we rely on the Gini index $\mathscr{G}$ [gini_measure_1936].
+Given a finite sequence of profits ranked in increasing order $J^{(1)}\leq J^{(2)}\leq\dots\leq J^{(n)}$, where $J^{(l)}$ denotes the $l$-th highest profit, $\mathscr{G}$ can be obtained as:
+$\mathscr{G}:=\frac{1}{n}\left(n+1-2\left(\frac{\sum_{i=1}^{n}(n+1-i)J^{(i)}}{\sum_{i=1}^{n}J^{(i)}}\right)\right).$
+The lower $\mathscr{G}$ is, the more equal is the sharing of profits.
 We report the results of this analysis in Tab.
@@
 Looking at the index computed by zone, the differences between classical and learning-based algorithm stay the same.
-TABLE IV: We report here the Gini index for all simulations: lower is a more equal share of the total profit between producers.
-Simulation Overall Germany Austria
-Best Response $0.43$ $0.16$ $0.15$
+TABLE IV: Gini index for all simulations.
+Simulation Overall $\mathscr{G}$ Germany $\mathscr{G}$ Austria $\mathscr{G}$
+Best-Response $0.43$ $0.16$ $0.15$
 Potential Solution $0.51$ $0.47$ $0.28$
@@
 In Tab.
-V we report the Evaluation and Training time: the two classical algorithms obviously have no training time, while for the MARL algorithm, the latter is what takes the longest.
-Although, it is important to stress that the MARL algorithm needs to be trained once and then it can be used to evaluate multiple settings without retraining, or eventually with a much shorter training routine.
-We also highlight that the BR algorithm’s evaluation time is longer than the training of 8 RL agents, due to the generally slow convergence.
+V we report the evaluation and training times: the two classical algorithms obviously have no training time, while this is the most time-consuming part for MARL.
+Further, MARL achieves faster evaluation time than the exact methods.
+The BR algorithm’s evaluation time is longer than the training of 8 RL agents, due to the sequential calls of the optimization solver.
 TABLE V: Algorithms time for evaluation and training.
 Algorithm Evaluation Time Training Time
-Best Response $\sim 8$h -
+Best-Response $\sim 8$h -
 Potential $\sim 10$m -
@@
 (6).
-We write the export limit for Germany (Austria) as a fraction of the maximal demand of Austria (Germany): $E_{G}=c_{G}\cdot\max{D_{A}}$, and we explore the ranges $c_{G}\in[0,2]$ and $c_{A}\in[0,0.2]$, so we go from completely uncoupled zone to a situation where the larger zone (Germany) has an export limit which is double the total demand of the smaller zone (Austria).
+We write the export limit for Germany (Austria) as a fraction of the maximal demand of Austria (Germany): $E_{G}=c_{G}\cdot\max{D_{A}}$, and we explore the ranges $c_{G}\in[0,2]$ and $c_{A}\in[0,0.2]$, thus we go from uncoupled zones to a scenario where the larger zone (Germany) has an export limit which is double the total demand of the smaller zone (Austria).
 To quickly evaluate all of these values we use the MARL simulations, in both cases (4 and 8 RL agents).
@@
 3, for the range of constraints values.
-In particular, we trained the agents for the two cases with $c_{G}=0.4,c_{A}=0.04$, and used them on all other values.
-When analyzing these results, the actual costs are not particularly relevant, as they depend on the demand and other external factors, while the trend of increase/decrease is what we have to interpret.
+In particular, we trained the agents with $c_{G}=0.4,c_{A}=0.04$, and used them on all other values.
 We can see that in the case of $4$ RL agents, the average cost is not influenced by the export constraint: this is likely due to the fact that the presence of static agents, always bidding at marginal price, already drives the cost down, and it is not possible to decrease it further.
@@
 Furthermore, since it is the market that decides how to divide each bid between zone, a producer can not create a priori a specific bid for a specific zone, each bid needs to be general enough to be competitive on every zone available.
-We conclude by mentioning [24], where the authors simulated a more economically sophisticated model of the European electricity market, to test what the addition of another smaller zone would do to the costs in the French zone.
-Their results show the same trend we observe here, with the smaller zone obtaining similar cost, while the larger one decreases its own.
-Our main methodological advantage, using a MARL based simulation, is that we can quickly test a range of different settings, not just a single scenario, providing a less detailed but broader view of the effect of different market designs.
-Further, using pre-trained agents, we are testing how a bidding strategy developed for a certain market design translates to a different one, when there is no time to adapt.
-This approach is of particular interest for situations where the coupling may change due to grid failures, with no warning.
+We conclude by mentioning [di_cosmo_welfare_2020], where the authors simulated a more economically sophisticated model of the European electricity market, to test what the addition of another smaller zone would do to the costs in the French zone.
+Their results show the same trend as ours, with the smaller zone obtaining similar cost, while the larger one decreases its own.
+The main advantage, using MARL based simulation, is that we can quickly test a range of different settings, not just a single scenario, providing a less detailed but broader view of the effect of different market designs.
+Further, using pretrained agents, we are testing how a bidding strategy developed for a certain market design translates to a different one, when there is no time to adapt.
+This is relevant for scenarios where the coupling may change due to grid failures, with no warning.
 VI Conclusions
-In this work we presented a model for multi-zonal ancillary markets based on a bilevel formulation.
-Each electricity producer is represented by an optimization problem with local and coupling constraints, with the objective of maximizing its profit, while the market operator objective is to minimize the total cost, so we create an adversarial formulation.
-We then reformulated the bilevel problem as a single level optimization, from the perspective of each producer, using the KKT reformulation of the market optimization.
-We interpret the ensemble of single level problems of each producer as a generalized Nash game.
-The game we obtain has side constraints, trilinear objective functions and nonconvex constraints, making it particularly challenging to solve.
-We analyze this game considering a penlaized version, and we obtain existence results for the Nash equilibrium that we relate to the initial generalized Nash game.
-We conclude our theoretical analysis showing that the generalized Nash game has a generalized potential structure.
-We then consider different algorithms specifically tailored to our problem: as baselines, we rely on an integrated optimization reformulation relying on the potential structure of the game, and Gauss-Seidel Best Response to solve the generalized Nash game.
-Further, we implement a multi-agent deep reinforcement learning algorithm to simulate the market in a more realistic, distributed way.
-We conclude validating the three algorithmic solutions on real data from the ancillary market between Germany and Austria.
-Our results showcase that computationally, the MARL simulation is the fastest to evaluate the data, but the agents have to be trained once before, while the best response approach is the slowest.
-The two classical algorithms also produce solutions with market costs much higher than the multi-agent counterpart, but the latter produces a more unfair split of the profits between the players.
-Finally, we test different forms of coupling between the zones, showing that a stronger coupling could lead to a decrease in cost for the larger zone, a result which is validated by other analysis present in the literature [24].
-Appendix A KKT Reformulation Details
-The full form of the Lagrangian for the market clearing problem of Eq. (8) is the following:
-$\begin{split}L(x;&\lambda,\mu,\sigma,\delta)=\sum_{n\in\mathcal{N}}\sum_{k=1}^{K_{n}}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\right)\Delta_{n,k}\pi_{n,k}\\ +&\sum_{z\in\mathcal{Z}}\lambda_{z}\left(D_{z}-\sum_{n\in\mathcal{N}}\sum_{k=1}^{K_{n}}x_{n,k}^{z}\Delta_{n,k}\right)\\ +&\sum_{n\in\mathcal{N}}\sum_{k=1}^{K_{n}}\mu_{n,k}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}-1\right)\\ +&\sum_{z\in\mathcal{Z}}\sigma_{z}\left(\sum_{n\in\mathcal{N}(z)}\sum_{k=1}^{K_{n}}\left(\sum_{z^{\prime}\neq z}x_{n,k}^{z^{\prime}}\right)\Delta_{n,k}-E_{z}\right)\\ +&\sum_{z\in\mathcal{Z}}\delta_{z}\left(C_{z}-\sum_{n\in\mathcal{N}(z)}x_{n,k}^{z}\Delta_{n,k}\right),\end{split}$ (18)
-where we explicit the dual variables $\lambda\in\mathbb{R}_{\geq 0}^{Z},\mu\in\mathbb{R}_{\geq 0}^{B},\sigma\in\mathbb{R}_{\geq 0}^{Z},\delta\in\mathbb{R}_{\geq 0}^{Z}$.
-Appendix B Penalized Objective Function
-We report here the explicit form for the penalized objective function used in the penalized GNG of Eqs. (14). As described in Section III-A, we relax constraints Eqs. (3)-(6), to obtain:
-$J_{n}^{p}(w)$ $=\sum_{1\leq k\leq K}\sum_{z\in\mathcal{Z}}\pi_{n,k}\Delta_{n,k}x_{n,k}^{z}$ (19a)
-$+\sum_{z\in\mathcal{Z}}M_{D_{z}}\left(\sum_{n\in\mathcal{N}}\sum_{1\leq k\leq K}x_{n,k}^{z}\Delta_{n,k}-D_{z}\right)$ (19b)
-$+M_{C}\left(\sum_{j\in z(n)}\sum_{0\leq k\leq K}x_{j,k}^{z(n)}\Delta_{j,k}-C_{z(n)}\right).$ (19c)
-Appendix C Proof of Proposition 1
-We call the two Lagrangian functions associated to the problems $\mathcal{L}_{n}^{P}$ and $\mathcal{L}_{n}^{U}$, and we consider the solution for $U$: $(w_{n}^{*},\Theta_{n}^{*})\in\mbox{SOL}\{\mbox{KKT of }U\}$, where $\Theta$ is the set of Lagrangian multipliers.
-This solution must respect all the constraints and be a solution of $\nabla\mathcal{L}^{U}(w^{*}_{n},\Theta^{*}_{n})=0$.
-We now call the part of Lagrangian multipliers that are still present in $P$ $\Theta[\Sigma_{n}]$, and we consider the set of equations coming from $\nabla\mathcal{L}^{P}(w_{n},\Theta_{n}[\Sigma_{n}];M_{n})=0$, where $M_{n}$ is the set of penalties for agent $n$.
-If we evaluate these equations in $(w_{n}^{*},\Theta_{n}^{*}[\Sigma_{n}])$, we can solve them w.r.t.
-the penalties $M_{n}$, obtaining $\bar{M}_{n}$.
-Given the assumptions, solving the system $\nabla\mathcal{L}^{P}(w_{n},\Theta_{n}[\Sigma];\bar{M}_{n})=0$ in $(w_{n},\Theta_{n}[\Sigma])$ provides $\tilde{w}_{n}$ which is a solution also for $\nabla\mathcal{L}^{U}(w_{n},\Theta_{n})=0$.
-Appendix D Hessians Analysis
-For the stationarity condition of Eq. (12c) we obtain the following cases:
-$x_{n}^{\top}\nabla_{x_{n}}^{2}\left(\nabla_{x_{n,k}^{z}}L\right)x_{n}=\begin{cases}\pi_{n,k}-\lambda_{n}^{z}-\delta_{n}\geq 0,\quad\mbox{if }z=z(n)\\ \pi_{n,k}-\lambda_{n}^{z}+\sigma_{n}\geq 0,\quad\mbox{if }z\neq z(n)\end{cases}$
-Given that an interior point forces $\lambda_{n}^{z}=0,\delta_{n}=0$ and $\sigma_{n}=0$, and that the prices $\pi_{n,k}$ are positive by definition, both inequalities are always satisfied.
-Regarding the Hessian coming from Eq.
-(5), which we call $\nabla^{2}G_{D_{z}}$, we can observe that deriving twice we get only constant terms, and given that all the terms in the sum are positive, we obtain a Hessian matrix with only $1$s.
-Then, since our decision variables are all positive by definition, the requirement $x^{\top}\nabla^{2}G_{D_{z}}x\geq 0$ is always true.
-Appendix E MADDPG Implementation Details
-There are three main problems related to the use of RL algorithms: defining the input state and the output action, mapping the action to satisfy the constraints and deciding the reward function.
-State and Action
-We choose the following forms for the input state and output action, for each producer:
-  • state: $s^{t}_{z}=[p^{t-1},d^{t-1},w^{t-1},\sigma^{t-1}_{z}]$ where $p^{t-1}$ are the previous day clearing prices, $d^{t-1}$ is the day, $w^{t-1}$ is the day of the week and $\sigma_{z}^{t-1}$ is a random signal which is zone-specific.
-  • action: $a^{t}_{i}=\left[\left[\pi_{n}^{t,0},\dots,\pi_{n}^{t,K_{n}}\right],\left[\Delta_{n}^{t,0},\dots,\Delta_{i}^{t,K_{n}}\right]\right]$, where $K_{n}$ is the number of bids each producer can submit.
-Action Mapping
-In our market model we have the following features:
-  • Capacity Output: we want each capacity bid to be at least 5 MW and less than the maximum capacity, with also the sum being less or equal than the total capacity of the producer.
-  • Price Output: each price bid must be larger than the marginal price $\underline{\pi}_{n}$ and smaller than the maximum bid $\bar{\pi}_{n}$
-We tackle the two outputs separately, building a module for the capacity and one for the price. We use the following mapping:
-  • Capacity Module: The output of the capacity module is a vector of length equal to the length of the bid sequence, which is passed through the soft-max function, representing the percentage of capacity to offer at each step.
-The benefit of this mapping is that is rather simple, with the drawback of forcing each producer to bid its whole capacity at each bid.
-  • Price Module: The price module produces two values $\alpha,\beta\geq 0$, which are used to obtain the lowest and highest price of the bidding sequence as: $\pi_{0}=\min(\bar{\pi},\underline{\pi}+\min(\alpha,\beta))$ and $\pi_{K}=\min(\bar{\pi},\underline{\pi}+\max(\alpha,\beta))$.
-We then compute the other prices interpolating between $\pi_{0}$ and $\pi_{K}$.
-Reward Design
-The most direct approach for the reward function in the market environment would be to use the profit, but it does not encode important information: a producer may place a non-optimal bid, but if the demand is high enough, it will be accepted anyway; further the zones have different demands, so a producer in a certain zone may have a smaller profit just because the demand is lower.
-What we propose is a reward formed by two terms: a normalized zone profit and a penalty for bid prices that are higher than the clearing price, weighted by the hyperparameter $\beta$:
-$r^{t}_{n}=\sum_{z\in\mathcal{Z}}\left(\frac{p_{n,z}^{t}}{d^{t}_{z}/|\mathcal{N}(z)|}-\beta\sum_{k=1}^{K_{n}}\mathbb{I}[\lambda_{z}^{t}>\underline{\pi}_{n}]\cdot(\pi_{n}^{t,k}-\lambda_{z}^{t})\right)$ (20)
+In this work, we rely on an adversarial Stackelberg game model of the multi-zonal ancillary markets, to study the impact of zonal coupling, that we recast as a nonconvex generalized Nash game with side constraints involving an associated price clearance to be satisfied by the equilibria.
+Further, relying on the potential structure of the generalized Nash game, we show the existence of a Nash equilibrium.
+Equilibrium can be reached making use of the game structure and relying on reformulations.
+Specifically, taking advantage of the potential structure of the game, we compare two exact approaches: Gauss-Seidel best-response and an integrated optimization reformulation.
+Another paradigm relies on the implementation of a multi-agent reinforcement learning (MARL) framework to simulate decentralized agent interactions, allowing us to study strategic behaviors and their impact on market dynamics.
+We validate the three algorithmic approaches using real data from the ancillary market between Germany and Austria.
+Our results show that, computationally, the MARL simulation (counting training and evaluation time) is more efficient than the best-response method, while remaining a decentralized approach.
+The two exact algorithms produce solutions with significantly higher total market costs compared to MARL; although the latter results in a more unfair allocation of profits among players.
+Numerically, we analyze different levels of coupling between zones, highlighting that tighter coupling may reduce costs in the larger zone; a result which is consistent with findings reported in the literature, e.g., in [di_cosmo_welfare_2020].